### v2

# DAOcc: 3D Object Detection Assisted Multi-Sensor Fusion for 3D Occupancy Prediction

Abstract
Multi-sensor fusion significantly enhances the accuracy and robustness of 3D semantic occupancy prediction, which is crucial for autonomous driving and robotics. However, most existing approaches depend on large image resolutions and complex networks to achieve top performance, hindering their application in practical scenarios. Additionally, most multi-sensor fusion approaches focus on improving fusion features while overlooking the exploration of supervision strategies for these features. To this end, we propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection supervision to assist in achieving superior performance, while using a deployment-friendly image feature extraction network and practical input image resolution. Furthermore, we introduce a BEV View Range Extension strategy to mitigate the adverse effects of reduced image resolution. Experimental results show that DAOcc achieves new state-of-the-art performance on the Occ3D-nuScenes and SurroundOcc benchmarks, and surpasses other methods by a significant margin while using only ResNet50 and 256×704 input image resolution. Code will be made available at https://github.com/AlphaPlusTT/DAOcc.
1 Introduction
3D semantic occupancy prediction (occ) is a critical task in autonomous driving [54, 4, 50] and robotic systems [45, 47, 15, 48], where accurately understanding the environment is essential for safe and efficient navigation. Reliable occupancy prediction requires not just accurate spatial data but also a comprehensive understanding of the environment’s context. Achieving this necessitates the integration of data from multiple sensors. LiDAR provides precise 3D spatial information for obstacle detection, while cameras capture visual details like color and texture for a deeper understanding of the scene. By combining these complementary data sources, the accuracy and robustness of occupancy prediction are significantly enhanced.
In existing works on multi-modal [60, 37, 61, 52, 43, 49, 38] or image-based [4, 51, 44, 13, 50, 24, 57, 42, 53] occ tasks, achieving superior performance often involves using extremely high image resolutions and complex image feature extraction networks (see Table 1 and Table 3), such as using 900×1600 resolution input image and ResNet101 [12] equipped with DCN [8, 66]. However, this approach significantly limits the deployment of top-performing occ models on edge devices due to the high computational demands. In comparison to images, point cloud is much sparser. For example, in the training set of the nuScenes [3] dataset, the maximum number of points in a single frame point cloud is only 34,880, which is equivalent to just 2.4% of the number of pixels in a 900×1600 resolution image. Therefore, how to effectively leverage point cloud data within a multi-modal occ framework remains to be further explored.
Moreover, we observe that most works [60, 37, 61, 52, 43, 49] on multi-modal occ primarily focuses on obtaining more effective fusion features, with insufficient exploration into the forms of supervision for these fused features. Although CO-Occ [38] introduces regularization based on implicit volumetric rendering to supervise fused features, it only utilizes distance ground truth from the original point cloud data, failing to fully exploit the geometry and structure information inherent in point cloud features. In contrast, LiDAR-based 3D detectors [59, 7, 34] effectively leverage this information, achieving significantly better performance in 3D object detection tasks compared to image-based 3D detectors [28, 17, 29, 30]. This observation suggests a new research direction: how to better exploit the unique strengths of point cloud data in multi-modal occ tasks.
Based on these observations, we propose DAOcc, a novel multi-modal occ framework that leverages 3D object detection to assist in achieving superior performance while using a deployment-friendly image encoder and practical input image resolution.
In constructing the baseline network of DAOcc, we adopt the simplest and most direct approach. First, we use a 2D image encoder and 3D sparse convolutions [55] to extract features from images and point clouds, respectively. Given that depth estimation from monocular images is an ill-posed problem [14, 21] and that deformable attention modules are overly complex [11], we employ a simple method similar to Harley et al. [10] to transform image features from 2D space to 3D volumetric space. Specifically, we project a set of predefined 3D points onto the 2D image feature plane and use bilinear interpolation to sample the corresponding 2D image features for these points. Next, we adopt a straightforward fusion strategy by concatenating the image and point cloud features and then applying 2D convolution to obtain a unified BEV (Bird’s Eye View) feature. Finally, we apply a fully convolutional BEV encoder with a residual structure to further fuse the unified BEV features, and subsequently employ the Channel-to-Height [57] operation to transform the channel dimension into the height dimension.
To fully exploit the geometry and structure information inherent in point cloud features, we augment the baseline model’s unified BEV features with 3D object detection supervision, thereby enhancing the discriminability of the unified BEV features. This renders the unified BEV features more sensitive to object boundaries and enables our method to avoid the need for complex loss function combinations, as depicted in Figure 1. Additionally, given the sparsity of point clouds, we extend the processing range of the point clouds and employ sparse convolutions [55] to mitigate the computational overhead introduced by this extension. We refer to this approach as BVRE (BEV View Range Extension). BVRE provides a larger BEV field of view, offering more contextual information and mitigating the adverse effects of reduced image resolution. It is worth noting that the 3D object detection supervision is only used as an auxiliary branch during training and can be removed during inference for the occ task.
Extensive experiments are conducted to validate the efficiency and effectiveness of the proposed DAOcc. The results show that DAOcc establishes the new state-of-the-art performance on the Occ3D-nuScenes [44] and SurroundOcc [51] benchmarks, while using only ResNet50 [12] and a 256x704 input image resolution. Specifically, on the Occ3D-nuScenes validation set, DAOcc achieves 53.82 mIoU when using the camera mask during training and 48.2 RayIoU without using the camera mask. Additionally, on the SurroundOcc validation set, DAOcc achieves 45.0 IoU and 30.5 mIoU.
In conclusion, our contributions are summarized as follows:
We design a simple yet efficient multi-modal 3D semantic occupancy prediction baseline, eliminating the need for complex deformable attention [23] modules as well as image depth estimation during feature fusion.
We propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection to assist in achieving superior performance while using a deployment-friendly image encoder and practical input image resolution.
We introduce a BEV View Range Extension strategy, which provides a larger BEV field of view, offering more contextual information and mitigating the adverse effects of reduced image resolution.
We establish the new state-of-the-art performance on Occ3D-nuScenes dataset and SurroundOcc dataset, while using only ResNet50 and a 256×704 input image resolution.
2 Related Work
3D Occupancy Prediction: 3D occupancy prediction aims to map all occupied voxels in the environment and assign semantic labels, thus providing more fine-grained perception results.
Camera-based 3D occupancy prediction has gained significant attention due to its cost-effectiveness. Early works [4, 16, 62] lack dense surround occupancy supervision, resulting in sparse occupancy predictions. To obtain dense occupancy prediction, OpenOccupancy [49], SurroundOcc [51] and Occ3D [44] developed methods for generating dense occupancy labels and established benchmarks on their respective proposed datasets. It is worth noting that SurroundOcc does not introduce manual annotation. The emergence of these benchmarks further promoted the development of 3D occupancy prediction. PanoOcc [50] adopts voxel query and a coarse-to-fine approach to learn a unified occupancy representation. FlashOcc [57] uses BEV features with 2D convolution to achieve efficient feature extraction and proposes the Channel-to-Heigh transformation to lift the output of BEV to 3D space. SparseOcc [31] leverages a sparse voxel decoder to reconstruct the sparse geometry of a scene and uses sparse queries to predict masks and labels. However, all camera-based methods inevitably need to cope with the challenges of adverse lighting and weather conditions, and many works inevitably lack robustness due to the ill-posed monocular depth estimation problem [14, 21].
While camera-based occupancy prediction has demonstrated promising results, multi-modal approaches offer greater reliability and robustness, making them indispensable for practical applications in autonomous driving and robotics. Due to the fact that cameras are susceptible to changing lighting and weather conditions, OccFusion [37] enhances the accuracy and robustness of the occupancy network by integrating features from LiDAR and radar. In concurrent work, OccFusion [61] projects the preprocessed denser and more uniform point cloud onto the image plane to establish a mapping relationship, and performs deformable attention [67] to fuse the corresponding features. Although OccFusion [61] avoids depth estimation, using deformable attention incurs a greater computational burden [11]. Hydra [52] extends FB-BEV [25] to the camera-radar fusion domain and improves the accuracy of depth estimation with the help of radar features. Co-Occ [38] employs K-nearest neighbor search in selecting neighboring camera features to enhance the corresponding LiDAR features and proposes a regularization based on implicit volume rendering. However, this regularization only utilizes the distance ground truth of the point cloud and does not leverage its intrinsic geometry information. EFFOcc [43] proposes an efficient and lightweight multi-modal 3D occupancy network, but it significantly relies on the 3D object detection pre-training of the LiDAR branch for optimal results. This dependence, however, limits the flexibility of the network structure design. Specifically, to maximize the benefits of the 3D detection pre-training, EFFOcc has to adjust its network architecture to align more closely with well-established LiDAR detection networks. In this work, we introduce a simple yet efficient multi-modal occupancy prediction network, eliminating the need for complex deformable attention [23] as well as image depth estimation during feature fusion, and does not require 3D object detection pre-training. Furthermore, to fully exploit the inherent geometry information in point cloud features, we add supervision of 3D object detection as an auxiliary supervision on the fused features.
Multi-Modal 3D Object Detection: Recent multi-modal 3D object detection methods [33, 19, 20, 6, 46] mainly focus on learning effective BEV feature representations. TransFusion [1] proposes a two-stage transformer-decoder based detection head and applies cross attention to obtain image features for each object query. BEVFusion [33] proposes an efficient and generic multi-task multi-sensor fusion framework which unifies multi-modal features in the shared BEV representation space, and introduces a specialized kernel to speed up the BEV pooling operation. In concurrent work, BEVFusion [26] decomposes the LiDAR-camera fusion into two streams that can independently output perception results, and performs feature fusion after the two streams. DAL [14] follows the concept of ‘Detecting As Labeling’ and decouples the fused features during classification and regression. Specifically, it employs the fused features for classification, whereas it utilizes the point cloud features exclusively for regression. In this work, we incorporate the simple yet effective feature fusion approach of BEVFusion [33] and introduce 3D object detection as an auxiliary branch during training.
3 Proposed Method
3.1 Overall Framework
Our purpose is to fully utilize the inherent geometry and structure information in point cloud features within a multi-modal occupancy prediction framework. Previous multi-modal methods have not fully exploited this and can only achieve superior performance through a more complex image encoder and a larger input image resolution. The overall framework of our proposed DAOcc is illustrated in Figure 2. DAOcc takes surround images and their corresponding time-synchronized point cloud as input, and obtains the features of the image and point cloud through the Camera Encoder and LiDAR Encoder, respectively. The 2D image features are transformed into the 3D voxel space by projection and interpolation. Subsequently, the image and point cloud features in the 3D space are compressed along the height dimension to generate the corresponding BEV(Bird’s Eye View) features. A simple 2D convolution is then applied for feature fusion, and the fully convolutional BEV Encoder encodes these fused features to obtain the final BEV representations. Finally, the occupancy head uses the Channel-to-Height operation [57] to restore the height of the BEV representations, resulting in final 3D voxel space representations that can be utilized for occupancy prediction. These modules collectively form the basic network architecture of DAOcc, which will be elaborated in Sec. 3.2.
On the foundation of the basic network, we introduce the BVRE strategy (Sec. 3.3) to compensate for the information loss that arises from the reduction in image resolution. This strategy aims to enrich the spatial contextual semantic information by broadening the perceptual range from the BEV perspective. Furthermore, to fully leverage the inherent geometry and structure information within point cloud features, we incorporate 3D object detection as auxiliary supervision (Sec. 3.4). This auxiliary supervision not only enhances the discriminability of the fused features but also leads to a very concise overall training loss (Sec. 3.5) for our proposed framework.
3.2 Basic Network
LiDAR Encoder. The method of embedding raw LiDAR points into 3D voxelized features is consistent with SECOND [55]. We first voxelize the point cloud, retaining a maximum of 10 points per voxel, resulting in a 3D voxel grid of size .The feature representation for each voxel is obtained by averaging the features of all points within it. Next, we apply 3D sparse convolutions [55] to encode these voxel features, generating spatially compressed LiDAR voxel features , where represents the feature dimensions.
Camera Encoder. For image feature extraction, taking surround images as input, we first use ResNet50 [12] as the backbone to extract multi-scale features, denoted as , where represent features extracted after downsampling. Then, we employ the Feature Pyramid Network (FPN) [27] as the neck to aggregate these multi-scale features. The output feature map has a shape of , where and represent the input resolution of the image, and and indicate the number of channels and the number of surround images, respectively.
Projection and Interpolation. For image-related occupancy prediction, a key step is to transform image features from 2D image plane to 3D volume space. Most existing methods use monocular depth estimation [49, 43, 13, 57] or deformable attention [51, 44, 50, 61]. However, monocular depth estimation is inherently an ill-posed problem [14, 21], while deformable attention imposes a significant computational burden [11]. Given these limitations, we use a simple yet effective projection and sample approach similar to Harley et al. [10]. Concretely, we first predefine a 3D voxel grid with a shape of , where represents the number of voxels along the -axis. The center point of each voxel is then projected onto the image feature plane, and only points that fall within both the image feature plane and the camera’s field of view are retained. Next, the sub-pixel projection positions of the retained points are bilinearly interpolated to generate the image features corresponding to each voxel. For voxels located in the overlapping viewing regions of the surround cameras, we average the image features from the two corresponding cameras to obtain the final feature for each voxel. The output camera’s voxel features can be denoted as .
BEV Encoder. Given the fused feature , we further refine by passing it through three blocks of ResNet18 [12], resulting in two feature maps, and , extracted at two scales from the first and last blocks respectively. Then, similar to FPN [27], we apply bilinear upsampling to and concatenate it with along the feature dimension. Finally, we fuse the features of different scales using a convolution block. The output refined BEV features can be denoted as .
3.3 BVRE
As shown in Figure 3, due to the sparsity of point clouds, the increase in computational cost resulting from expanding the processing range of the point cloud is very small compared to increasing the image resolution. Therefore, we extend the point cloud range to provide more 3D spatial context, compensating for the information loss caused by lower image resolution. However, arbitrarily setting the range of the point cloud or voxel resolution may result in misalignment between the BEV features and the occupancy ground truth. Assuming that the voxel resolution of the occupancy grid is , the expanded point cloud data in BEV covers a rectangular region with ranging from to and ranging from to , and has a voxel resolution of . Thus, and must be integer multiples of , and must also be an integer multiple of . Otherwise, the spatial resolution represented by each feature vector along the -axis in the BEV representation will not be equal to . To avoid complex manual design, we adopt coordinate transformation and interpolation, as illustrated in the purple part of Figure 2, which can be formulated as follows:
where is a set of predefined points in the plane of the occupancy annotation coordinate system, each point is located at the center of an occupancy grid in the plane. is the transformation matrix from occupancy coordinates to LiDAR coordinates. The function of is to scale the coordinate values to a range from -1 to 1.
3.4 Detachable Auxiliary Detection Head
To fully exploit the inherent geometry and structure information in point cloud features, we employ a detachable auxiliary detection head to further supervise the feature fusion process, thereby enhancing the information about object boundaries and scene structures in the fused features. Meanwhile, the auxiliary detection task is related to the occupancy prediction task, thus providing multiple regularization effects during optimization. For simplicity, we use the one-stage CenterHead introduced in CenterPoint [56] as the auxiliary detection head. Given the refined BEV feature , we utilize two convolutional layers to generate a keypoint heatmap from the BEV perspective, and apply a Gaussian kernel [18, 64] to map the center points of all ground truth 3D boxes onto a target heatmap . The training objective is formulated as a focal loss [41] based on Gaussian heatmaps:
where and represent the predicted score and ground truth of the heatmap at location , respectively. is the number of objects in a point cloud, and and are hyper-parameters of the focal loss [41]. Similarly, for the offset of the center point, as well as the height-above-ground, 3D dimensions and yaw of the 3D bounding box, we use separate convolutional layers to predict each of these parameters, and then apply the L1 loss for supervision:
where and represent the predicted value and the ground truth respectively. The above two tasks enable the network to perceive object boundaries more precisely, which in turn facilitates achieving more precise occupancy predictions. The total loss for the auxiliary detection head can be denoted as:
3.5 Overall Objective Function
Most existing methods rely on complex combinations of loss functions [49, 50, 24, 58, 42, 53, 37, 61, 38, 43], such as a combination of several or all of the focal loss [41], the scene-class affinity loss [4], the dice loss, the lovasz-softmax loss [2], the depth loss [22], etc., to achieve the expected performance. In contrast, our approach utilizes a detachable 3D object detection branch for auxiliary supervision and requires only a simple cross-entropy loss for occupancy prediction. The total loss for our framework can be defined as:
4 Experiments
Method | mIoU | Modality | Image Size | 2D Backbone |
others |
barrier |
bicycle |
bus |
car |
const. veh. |
motorcycle |
pedestrian |
traffic cone |
trailer |
truck |
drive. suf. |
other flat |
sidewalk |
terrain |
manmade |
vegetation |
BEVDetOcc(2f) [13] | 42.02 | C | 512 1408 | Swin-B | 12.2 | 49.6 | 25.1 | 52.0 | 54.5 | 27.9 | 28.0 | 28.9 | 27.2 | 36.4 | 42.2 | 82.3 | 43.3 | 54.6 | 57.9 | 48.6 | 43.6 |
PanoOcc(4f) [50] | 42.13 | C | 864 1600 | R101-DCN | 11.7 | 50.5 | 29.6 | 49.4 | 55.5 | 23.3 | 33.3 | 30.6 | 31.0 | 34.4 | 42.6 | 83.3 | 44.2 | 54.4 | 56.0 | 45.9 | 40.4 |
FB-OCC(16f) [24] | 48.90 | C | 960 1760 | VoVNet-99 | 14.3 | 57.0 | 38.3 | 57.7 | 62.1 | 34.4 | 39.4 | 38.8 | 39.4 | 42.9 | 50.0 | 86.0 | 50.2 | 60.1 | 62.5 | 52.4 | 45.7 |
FlashOcc(2f) [57] | 43.52 | C | 512 1408 | Swin-B | 13.4 | 51.1 | 27.7 | 51.6 | 56.2 | 27.3 | 30.0 | 29.9 | 29.8 | 37.8 | 43.5 | 83.8 | 46.6 | 56.2 | 59.6 | 50.8 | 44.7 |
COTR(8f) [36] | 46.21 | C | 512 1408 | Swin-B | 14.9 | 53.3 | 35.2 | 50.8 | 57.3 | 35.4 | 34.1 | 33.5 | 37.1 | 39.0 | 45.0 | 84.5 | 48.7 | 57.6 | 61.1 | 51.6 | 46.7 |
RadOcc [60] | 49.38 | C+L | 512 1408 | Swin-B | 10.9 | 58.2 | 25.0 | 57.9 | 62.9 | 34.0 | 33.5 | 50.1 | 32.1 | 48.9 | 52.1 | 82.9 | 42.7 | 55.3 | 58.3 | 68.6 | 66.0 |
OccFusion [37] | 46.67 | C+L+R | 900 1600 | R101-DCN | 12.4 | 50.3 | 31.5 | 57.6 | 58.8 | 34.0 | 41.0 | 47.2 | 29.7 | 42.0 | 48.0 | 78.4 | 35.7 | 47.3 | 52.7 | 63.5 | 63.3 |
OccFusion [61] | 48.74 | C+L | 900 1600 | R101 | 12.4 | 51.8 | 33.0 | 54.6 | 57.7 | 34.0 | 43.0 | 48.4 | 35.5 | 41.2 | 48.6 | 83.0 | 44.7 | 57.1 | 60.0 | 62.5 | 61.3 |
HyDRa [52] | 44.4 | C+R | 256 704 | R50 | - | - | - | 52.3 | 56.3 | - | 35.9 | 35.1 | - | - | 44.1 | - | - | - | - | - | - |
EFFOcc [43] | 52.62 | C+L | 512 1408 | Swin-B | 14.7 | 58.6 | 34.7 | 60.5 | 65.5 | 36.0 | 40.4 | 51.5 | 41.0 | 48.5 | 54.6 | 84.2 | 46.4 | 57.9 | 60.8 | 70.9 | 68.4 |
DAOcc (Ours) | 53.82 | C+L | 256 704 | R50 | 12.4 | 59.6 | 38.4 | 61.9 | 67.1 | 35.3 | 48.2 | 59.1 | 43.5 | 50.9 | 56.3 | 83.0 | 44.7 | 56.7 | 59.9 | 70.0 | 68.1 |
Method | RayIoU | Modality | Image Size | 2D Backbone | RayIoU1m | RayIoU2m | RayIoU4m | mIoU |
BEVFormer(4f) [23] | 32.4 | C | 900 1600 | R101 | 26.1 | 32.9 | 38.0 | 39.2 |
RenderOcc [39] | 19.5 | C | 512 1408 | Swin-B | 13.4 | 19.6 | 25.5 | 24.4 |
SimpleOcc [9] | 22.5 | C | 336 672 | R101 | 17.0 | 22.7 | 27.9 | 31.8 |
BEVDetOcc(8f) [13] | 32.6 | C | 384 704 | R50 | 26.6 | 33.1 | 38.2 | 39.3 |
FB-Occ(16f) [24] | 33.5 | C | 256 704 | R50 | 26.7 | 34.1 | 39.7 | 39.1 |
SparseOcc(16f) [31] | 36.1 | C | 256 704 | R50 | 30.2 | 36.8 | 41.2 | 30.9 |
Panoptic-FlashOcc [58] | 35.2 | C | 256 704 | R50 | 29.4 | 36.0 | 40.1 | 29.4 |
Panoptic-FlashOcc(8f) [58] | 38.5 | C | 256 704 | R50 | 32.8 | 39.3 | 43.4 | 31.6 |
DAOcc (Ours) | 48.2 | C+L | 256 704 | R50 | 44.4 | 48.7 | 51.6 | 47.9 |
Method | IoU | mIoU | Modality | Image Size | 2D Backbone |
barrier |
bicycle |
bus |
car |
const. veh. |
motorcycle |
pedestrian |
traffic cone |
trailer |
truck |
drive. suf. |
other flat |
sidewalk |
terrain |
manmade |
vegetation |
SurroundOcc [51] | 31.4 | 20.3 | C | 900 1600 | R101-DCN | 20.5 | 11.6 | 28.1 | 30.8 | 10.7 | 15.1 | 14.0 | 12.0 | 14.3 | 22.2 | 37.2 | 23.7 | 24.4 | 22.7 | 14.8 | 21.8 |
OccFormer [62] | 29.9 | 20.1 | C | 896 1600 | R101 | 21.1 | 11.3 | 28.2 | 30.3 | 10.6 | 15.7 | 14.4 | 11.2 | 14.0 | 22.6 | 37.3 | 22.4 | 24.9 | 23.5 | 15.2 | 21.1 |
C-CONet [49] | 26.1 | 18.4 | C | 896 1600 | R101 | 18.6 | 10.0 | 26.4 | 27.4 | 8.6 | 15.7 | 13.3 | 9.7 | 10.9 | 20.2 | 33.0 | 20.7 | 21.4 | 21.8 | 14.7 | 21.3 |
FB-Occ [24] | 31.5 | 19.6 | C | 896 1600 | R101 | 20.6 | 11.3 | 26.9 | 29.8 | 10.4 | 13.6 | 13.7 | 11.4 | 11.5 | 20.6 | 38.2 | 21.5 | 24.6 | 22.7 | 14.8 | 21.6 |
RenderOcc [39] | 29.2 | 19.0 | C | 896 1600 | R101 | 19.7 | 11.2 | 28.1 | 28.2 | 9.8 | 14.7 | 11.8 | 11.9 | 13.1 | 20.1 | 33.2 | 21.3 | 22.6 | 22.3 | 15.3 | 20.9 |
LMSCNet [40] | 36.6 | 14.9 | L | - | - | 13.1 | 4.5 | 14.7 | 22.1 | 12.6 | 4.2 | 7.2 | 7.1 | 12.2 | 11.5 | 26.3 | 14.3 | 21.1 | 15.2 | 18.5 | 34.2 |
L-CONet [49] | 39.4 | 17.7 | L | - | - | 19.2 | 4.0 | 15.1 | 26.9 | 6.2 | 3.8 | 6.8 | 6.0 | 14.1 | 13.1 | 39.7 | 19.1 | 24.0 | 23.9 | 25.1 | 35.7 |
M-CONet [49] | 39.2 | 24.7 | C+L | 896 1600 | R101 | 24.8 | 13.0 | 31.6 | 34.8 | 14.6 | 18.0 | 20.0 | 14.7 | 20.0 | 26.6 | 39.2 | 22.8 | 26.1 | 26.0 | 26.0 | 37.1 |
Co-Occ [38] | 41.1 | 27.1 | C+L | 896 1600 | R101 | 28.1 | 16.1 | 34.0 | 37.2 | 17.0 | 21.6 | 20.8 | 15.9 | 21.9 | 28.7 | 42.3 | 25.4 | 29.1 | 28.6 | 28.2 | 38.0 |
OccFusion [37] | 44.7 | 27.3 | C+L+R | 900 1600 | R101-DCN | 27.1 | 19.6 | 33.7 | 36.2 | 21.7 | 24.8 | 25.3 | 16.3 | 21.8 | 30.0 | 39.5 | 19.9 | 24.9 | 26.5 | 28.9 | 40.4 |
DAOcc (Ours) | 45.0 | 30.5 | C+L | 256 704 | R50 | 30.8 | 19.5 | 34.0 | 38.8 | 25.0 | 27.7 | 29.9 | 22.5 | 23.2 | 31.6 | 41.0 | 25.9 | 29.4 | 29.9 | 35.2 | 43.5 |
4.1 Datasets and Evaluation Metrics
Occ3D-nuScenes [44] is a large-scale benchmark dataset for 3D occupancy prediction, which contains 700 scenes for training and 150 scenes for validation. Each scene lasts for 20 seconds, with annotations provided at a frequency of 2 Hz. The dataset covers a perception range of [-40m, 40m] in the and directions, and [-1m, 5.4m] along the -axis, all within the ego coordinate system. This range is discretized into voxels, each with a size of [0.4m, 0.4m, 0.4m]. Each occupied voxel is assigned one of 17 semantic labels, including 16 common classes and a general object class labeled as ‘others’. The dataset includes RGB images captured by six surrounding cameras and LiDAR point clouds, enabling dense voxel-wise annotations for 3D scene understanding. Additionally, the dataset provides visibility masks for both LiDAR and camera modes, which can be utilized for training purposes.
SurroundOcc [51] is an automatically generated occupancy prediction dataset that does not require extra human annotation. It is built upon the nuScenes [3] and generates dense occupancy ground truth by utilizing existing 3D detection and 3D semantic segmentation labels. The dataset uses the nuScenes training set to train the model and employs the nuScenes validation set for evaluation purposes. The occupancy prediction range is set to [-50m, 50m] for the and axes, and [-5m, 3m] for the -axis, all within the LiDAR coordinate system. The final output 3D occupancy grid has a shape of 200×200×16, with each voxel measuring [0.5m, 0.5m, 0.5m].
Evaluation metrics: Following previous methods, we use the mean Intersection over Union (mIoU) as our evaluation metric when training models with camera visible mask provided by Occ3D-nuScenes, ensuring a fair comparison with existing approaches. When the camera visible mask is excluded from the training phase, we report an additional semantic segmentation metric, RayIoU [31], which has been recently introduced. For the SurroundOcc dataset, we report semantic segmentation performance using IoU and mIoU. Here, IoU is defined as the intersection over union of occupied voxels, calculated while ignoring their semantic categories, and it is used as an evaluation metric for the scene completion (SC) task.
4.2 Implementation Details
We use ResNet-50 [12], pre-trained on nuImages [3] by MMDetection [5], as our image backbone and adjust the input image resolution to 256×704. In the LiDAR branch, we aggregate LiDAR sweeps from 10 adjacent timestamps as input and initiate training from scratch. Data augmentation techniques, such as scaling, rotation, and flipping, are applied to the input images, and we also perform random flipping in the BEV space. The loss weights are configured as and . We adopt the AdamW optimizer [35] with a cosine annealing learning rate scheduler, incorporating a warmup phase, and set the initial learning rate to 2e-4. Unless conducting ablation experiments, we train our DAOcc using CBGS [65] for 6 epochs, the input point cloud range is set to [-54.0, -54.0, -5.0, 54.0, 54.0, 3.0] and the voxel size is set to [0.075, 0.075, 0.2]. All models are trained on 8 NVIDIA RTX4090 GPUs with a batch size of 4. Additionally, if the camera visible mask is not used during training, we randomly discard 80% of empty voxels, consistent with Zhao et al. [63]. Temporal information fusion, exponential moving average (EMA) for weight parameter and test-time augmentation is not used.
4.3 Comparison with State-of-the-Art Methods
Method | Modality | 3D Backbone | Param | mIoU |
FB-Occ [24] | C | - | 130.8M | 48.90 |
RadOcc [60] | C+L | SECOND | 135.4M | 49.38 |
EFFOcc [43] | C+L | SECOND | 108.6M | 52.62 |
DAOcc | C+L | SECOND | 56.2M | 53.82 |
Comparison on Occ3D-nuScenes with camera mask. As presented in Table 1, we compare our proposed DAOcc with both image-based and multi-modal methods on the Occ3D-nuScenes dataset. All results are provided either directly by their authors or based on their official code implementations. Our DAOcc establishes a new state-of-the-art performance, achieving an outstanding 53.82 mIoU, while utilizing a 2D backbone ResNet50 [12] and 256×704 input image resolution. Compared with the latest state-of-the-art multi-modal method EFFOcc [43], DAOcc achieves improvement of 1.2 mIoU. Notably, the image feature extraction network Swin-B [32] in EFFOcc has approximately 3 times the number of parameters of ResNet50, and EFFOcc uses a larger input image resolution of 512×1408. This highlights our method’s ability to mine more useful information from point cloud data. In addition, to achieve optimal results, EFFOcc requires the LiDAR branch to undergo 3D object detection pre-training, which leads to an improvement of approximately 3 mIoU. In contrast, DAOcc does not require LiDAR branch pre-training, but instead seamlessly integrates 3D object detection as an auxiliary branch into the multi-modal occupancy prediction framework, simplifying the model training process and achieving better performance.
Furthermore, we compare the complexity of the models, as shown in Table 4. Compared to the two top-performing multi-modal methods, RadOcc [60] and EFFOcc [43], as well as the top-performing image-based method FB-Occ [24], our proposed DAOcc has only about half the number of parameters but achieves significantly better performance (outperforming RadOcc, EFFOcc, and FB-Occ by 4.44, 1.2 and 4.92 mIoU).
Comparison on Occ3D-nuScenes without camera mask. To the best of our knowledge, no existing multi-modal occupancy prediction work has used RayIoU metrics to report model performance, so we mainly compare with image-based methods. As shown in Table 2, benefiting from the effective use of multi-modal input, especially point cloud data, our DAOcc achieves an outstanding 48.2 RayIoU and significantly outperforms image-based methods by a large margin, even though many of these methods use multi-frame temporal fusion. It should be noted that some image-based methods, such as Panoptic-FlashOcc and FB-Occ [24], adopt the Explicit Depth Supervision [22] and leverage depth information derived from point cloud data to supervise the training of monocular depth estimation from images, so these methods can also be regarded as multi-modal during the training phase.
Comparison on SurroundOcc. Table 3 presents a quantitative comparison on the SurroundOcc validation set, showcasing the performance of our DAOcc against other methods. Leveraging both camera and LiDAR inputs, DAOcc achieves a notable 3.2 mIoU improvement over OccFusion [37], which incorporates radar data in addition to camera and LiDAR inputs. It is worth noting that OccFusion utilizes R101-DCN with a high-resolution input of 900×1600, whereas our method employs a more lightweight Resnet50 backbone and a lower input resolution of 256×704. This underscores the effectiveness and efficiency of our proposed method.
4.4 Ablation Studies
Ablation studies are conducted on the validation set of Occ3D-nuScenes, with the camera visible mask being used during training. Unless otherwise specified, all experiments presented in this section are trained for 15 epochs without employing CBGS [65].
BEV View Range (m) | BEV Voxel Size (m) | mIoU | |
(a) | 0.075 0.075 | 50.76 | |
(b) | 0.075 0.075 | 50.98 | |
(c) | 0.1 0.1 | 49.75 | |
(d) | 0.1 0.1 | 49.98 | |
(e) | 0.075 0.075 | 51.26 |
Effectiveness of the BVRE. We propose the BVRE strategy to enrich spatial contextual information by expanding the perceptual range from a BEV (Bird’s Eye View) perspective. As shown in Table 5, a comparison of (a) and (b) indicates that, with the same voxel size, expanding the BEV range from [-41.4m, 41.4m] to [-45.6m, 45.6m] yields a performance improvement of 0.22 mIoU. Furthermore, the comparison of (a) and (e) shows that further expanding the BEV range to [-54.0m, 54.0m] augments the performance gain to 0.5 mIoU. This observation validates the value of spatial context information in a larger BEV range.
In addition, by comparing (b) and (c), we can find that the size of the voxel has a great impact on the model’s performance. To achieve better performance, we adopt the setting (e) in DAOcc.
camera mask | ADH | mIoU |
bicycle |
motorcycle |
pedestrian |
traffic cone |
✓ | 51.26 | 20.60 | 38.89 | 55.44 | 36.65 | |
✓ | ✓ | 52.82 (+1.56) | 36.35 (+15.75) | 45.48 (+6.59) | 58.27 (+2.83) | 41.98 (+5.33) |
44.31 | 14.94 | 32.22 | 43.77 | 34.07 | ||
✓ | 46.89 (+2.58) | 32.51 (+17.57) | 42.89 (+10.67) | 53.41 (+9.64) | 40.55 (+6.48) |
Ablation on the Auxiliary Detection Head. To fully leverage the inherent geometry and structure information within point cloud features, we incorporate 3D object detection as auxiliary supervision. As shown in Table 6, by adding auxiliary 3D object detection supervision, we can achieve an improvement of 1.56 mIoU when using camera visible masks during training and 2.58 mIoU when not using the mask, respectively. This demonstrates that adding a detection branch can significantly enhance the discriminability of the multi-modal features. Furthermore, we report on the four categories that benefit the most from the 3D object detection supervision. Notably, all four categories are foreground objects, which further validates the effectiveness and rationale behind the performance improvement achieved by our method. Specifically, when the camera visible mask is used during training, the addition of the auxiliary detection branch results in improvements of 15.75 mIoU, 6.59 mIoU, 2.83 mIoU, and 5.33 mIoU for bicycles, motorcycles, pedestrians, and traffic cones, respectively. Conversely, when the camera visible mask is not used during training, the improvements are 17.57 mIoU, 10.67 mIoU, 9.64 mIoU, and 6.48 mIoU for the same categories, respectively.
5 Conclusion
In this paper, we propose a novel multi-modal occupancy prediction framework, DAOcc, which aims to achieve superior performance while utilizing a deployment-friendly image backbone and practical input image resolutions. DAOcc achieves this superior performance by fully leveraging the unique advantages of point cloud data within a multi-modal framework, advantages that have been largely overlooked by previous methods. Extensive experiments conducted on the Occ3D-nuScenes and SurroundOcc datasets demonstrate the superiority of DAOcc compared to existing methods. We believe that DAOcc has practical application potential and could inspire future research in multi-modal occupancy prediction.
Supplementary Material
Method | Modality | GPU Type | FPS | mIoU |
FlashOcc(2f) [57] | C | A100-PCIE-40G | 0.9 | 43.52 |
RadOcc [60] | C+L | - | 0.3 | 49.38 |
EFFOcc [43] | C+L | RTX A6000 | 1.8 | 52.62 |
DAOcc(ours) | C+L | RTX 3070Ti | 2.7 | 53.82 |
6 Runtime Comparisons
We compare the runtime of different methods, as shown in Table 7. The FPS is obtained by testing on the corresponding GPU Type listed in the table. Since the source code for RadOcc is unavailable and EFFOcc has not yet released its implementation, we rely on the runtime speeds reported in their respective papers as benchmarks. It’s worth noting that the computing capability of the RTX 3070 Ti is significantly lower than that of both the RTX A6000 and the A100. The results demonstrate that our proposed DAOcc achieves the best performance while maintaining the fastest runtime speed.
7 Ablation for Achieving Final Performance
In addition to the ablation studies presented in Section 4.4, this section provides further ablation experiments. Table 8 presents the ablation experiments that contributed to the final performance. First, by extending the training time from 15 epochs to 24 epochs, we can observe an improvement of 0.57 mIoU. Next, the use of CBGS [65] yields an additional 0.43 mIoU improvement. Since CBGS involves resampling, which increases the number of samples per epoch, we limit the training to 6 epochs to maintain a comparable total training time to the 24 epochs without CBGS for a fair comparison.
8 Detection Performance
Since our proposed DAOcc incorporates an auxiliary 3D object detection branch into the occupancy prediction framework, we also evaluate the performance of this auxiliary branch. The auxiliary detection branch is designed to identify 10 object categories, including cars, trucks, buses, trailers, construction vehicles, pedestrians, motorcycles, bicycles, traffic cones, and barriers. All experiments are conducted using the nuScenes [3] dataset for the auxiliary 3D object detection training, and the results are evaluated using the mean average precision (mAP) and nuScenes detection score (NDS) metrics provided by the nuScenes benchmark.
As shown in Table 9, when DAOcc is trained on the Occ3D-nuScenes [44] dataset (for occupancy prediction) and the camera visible mask is used during training, the auxiliary detection branch achieves a performance of 59.4 mAP and 64.3 NDS. Notably, comparable performance is observed when the camera visible mask is not used. However, when the SurroundOcc [51] dataset is employed for occupancy prediction training, the auxiliary detection branch achieves a lower performance of 53.6 mAP and 60.2 NDS. We attribute this performance gap to the fact that the surroundocc dataset is automatically generated without manual calibration, potentially introducing more noise and negatively affecting model convergence.
EP15 | EP24 | CBGS (EP6) | mIoU |
✓ | 52.82 | ||
✓ | 53.39 | ||
✓ | 53.82 |
9 Visualization
Figure 4 presents the visualizations of DAOcc on Occ3D-nuScenes validation set. The camera visible mask is not used during model training and visualization. The results demonstrate that DAOcc is capable of making a relatively complete prediction of the scene, accurately reproducing it in fine detail.
Dataset | Camera Mask | mIoU | mAP | NDS | Car | Truck | Bus | T.L. | C.V. | Ped. | M.C. | B.C. | T.C. | B.R. |
Occ3D-nuScenes [44] | ✓ | 53.8 | 59.4 | 64.3 | 85.7 | 57.2 | 65.8 | 33.1 | 22.8 | 75.5 | 68.3 | 47.3 | 70.2 | 68.2 |
Occ3D-nuScenes [44] | ✗ | 47.9 | 59.6 | 64.2 | 85.4 | 55.0 | 64.6 | 34.1 | 23.1 | 76.3 | 67.0 | 50.7 | 70.8 | 69.0 |
SurroundOcc [51] | - | 30.5 | 53.6 | 60.2 | 81.9 | 42.7 | 56.0 | 26.3 | 19.7 | 72.6 | 57.7 | 43.8 | 66.7 | 68.1 |

### v3

@@
 Multi-sensor fusion significantly enhances the accuracy and robustness of 3D semantic occupancy prediction, which is crucial for autonomous driving and robotics.
-However, most existing approaches depend on large image resolutions and complex networks to achieve top performance, hindering their application in practical scenarios.
-Additionally, most multi-sensor fusion approaches focus on improving fusion features while overlooking the exploration of supervision strategies for these features.
-To this end, we propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection supervision to assist in achieving superior performance, while using a deployment-friendly image feature extraction network and practical input image resolution.
-Furthermore, we introduce a BEV View Range Extension strategy to mitigate the adverse effects of reduced image resolution.
-Experimental results show that DAOcc achieves new state-of-the-art performance on the Occ3D-nuScenes and SurroundOcc benchmarks, and surpasses other methods by a significant margin while using only ResNet50 and 256×704 input image resolution.
-Code will be made available at https://github.com/AlphaPlusTT/DAOcc.
-1 Introduction
-3D semantic occupancy prediction (occ) is a critical task in autonomous driving [54, 4, 50] and robotic systems [45, 47, 15, 48], where accurately understanding the environment is essential for safe and efficient navigation.
+However, most existing approaches depend on high-resolution images and complex networks to achieve top performance, hindering their deployment in practical scenarios.
+Moreover, current multi-sensor fusion approaches mainly focus on improving feature fusion while largely neglecting effective supervision strategies for those features.
+To address these issues, we propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection supervision to assist in achieving superior performance, while using a deployment-friendly image backbone and practical input resolution.
+In addition, we introduce a BEV View Range Extension strategy to mitigate performance degradation caused by lower image resolution.
+Extensive experiments demonstrate that DAOcc achieves new state-of-the-art results on both the Occ3D-nuScenes and Occ3D-Waymo benchmarks, and outperforms previous state-of-the-art methods by a significant margin using only a ResNet-50 backbone and 256×704 input resolution.
+With TensorRT optimization, DAOcc reaches 104.9 FPS while maintaining 54.2 mIoU on an NVIDIA RTX 4090 GPU.
+Code is available at https://github.com/AlphaPlusTT/DAOcc.
+I Introduction
+3D semantic occupancy prediction (occ) is a critical task in autonomous driving [1, 2, 3] and robotic systems [4, 5, 6], where accurately understanding the environment is essential for safe and efficient navigation.
 Reliable occupancy prediction requires not just accurate spatial data but also a comprehensive understanding of the environment’s context.
@@
 By combining these complementary data sources, the accuracy and robustness of occupancy prediction are significantly enhanced.
-In existing works on multi-modal [60, 37, 61, 52, 43, 49, 38] or image-based [4, 51, 44, 13, 50, 24, 57, 42, 53] occ tasks, achieving superior performance often involves using extremely high image resolutions and complex image feature extraction networks (see Table 1 and Table 3), such as using 900×1600 resolution input image and ResNet101 [12] equipped with DCN [8, 66].
+In existing works on multi-modal [9, 10, 11, 12, 13] or image-based [2, 14, 15, 16, 3, 17, 18, 19, 20] occ tasks, achieving superior performance often involves using extremely high image resolutions and complex image feature extraction networks (see Table I), such as using 900×1600 resolution input image and ResNet101 [21] equipped with DCN [22, 23].
 However, this approach significantly limits the deployment of top-performing occ models on edge devices due to the high computational demands.
-In comparison to images, point cloud is much sparser.
-For example, in the training set of the nuScenes [3] dataset, the maximum number of points in a single frame point cloud is only 34,880, which is equivalent to just 2.4% of the number of pixels in a 900×1600 resolution image.
+In comparison to images, the point cloud is much sparser.
+For example, in the training set of the nuScenes [24] dataset, the maximum number of points in a single point cloud frame is only 34,880, which is equivalent to just 2.4% of the number of pixels in a 900×1600 resolution image.
 Therefore, how to effectively leverage point cloud data within a multi-modal occ framework remains to be further explored.
-Moreover, we observe that most works [60, 37, 61, 52, 43, 49] on multi-modal occ primarily focuses on obtaining more effective fusion features, with insufficient exploration into the forms of supervision for these fused features.
-Although CO-Occ [38] introduces regularization based on implicit volumetric rendering to supervise fused features, it only utilizes distance ground truth from the original point cloud data, failing to fully exploit the geometry and structure information inherent in point cloud features.
-In contrast, LiDAR-based 3D detectors [59, 7, 34] effectively leverage this information, achieving significantly better performance in 3D object detection tasks compared to image-based 3D detectors [28, 17, 29, 30].
+Moreover, we observe that most works [9, 10, 11, 12] on multi-modal occ primarily focuses on obtaining more effective fusion features, with insufficient exploration into the forms of supervision for these fused features.
+Although CO-Occ [13] introduces a regularization based on implicit volumetric rendering to supervise fused features, it only utilizes distance ground truth from the original point cloud data, failing to fully exploit the geometry and structure information inherent in point cloud features.
+In contrast, LiDAR-based 3D detectors [25, 26, 27] effectively leverage this information, achieving significantly better performance in 3D object detection tasks compared to image-based 3D detectors [28, 29, 30].
 This observation suggests a new research direction: how to better exploit the unique strengths of point cloud data in multi-modal occ tasks.
@@
 Finally, we apply a fully convolutional BEV encoder with a residual structure to further fuse the unified BEV features, and subsequently employ the Channel-to-Height [18] operation to transform the channel dimension into the height dimension.
-To fully exploit the geometry and structure information inherent in point cloud features, we augment the baseline model’s unified BEV features with 3D object detection supervision, thereby enhancing the discriminability of the unified BEV features.
-This renders the unified BEV features more sensitive to object boundaries and enables our method to avoid the need for complex loss function combinations, as depicted in Figure 1.
+To fully exploit the geometry and structure information inherent in point cloud features, we enhance the baseline model’s unified BEV features by incorporating 3D object detection supervision, thereby improving the discriminability of the unified BEV features.
+The detection-to-occupancy supervision injects explicit object-localization cues into the occupancy pipeline, sharpening object boundaries and boosting foreground class performance while keeping inference complexity low—since the detection branch is used only during training and removed at inference.
+It also eliminates the need for complex loss function combinations, as depicted in Figure 2.
 Additionally, given the sparsity of point clouds, we extend the processing range of the point clouds and employ sparse convolutions [31] to mitigate the computational overhead introduced by this extension.
+To address the misalignment between the expanded point cloud range and the occupancy label, we apply a combination of coordinate system transformation and feature interpolation to extract fusion features that are spatially aligned with the occupancy label.
 We refer to this approach as BVRE (BEV View Range Extension).
 BVRE provides a larger BEV field of view, offering more contextual information and mitigating the adverse effects of reduced image resolution.
-It is worth noting that the 3D object detection supervision is only used as an auxiliary branch during training and can be removed during inference for the occ task.
 Extensive experiments are conducted to validate the efficiency and effectiveness of the proposed DAOcc.
-The results show that DAOcc establishes the new state-of-the-art performance on the Occ3D-nuScenes [44] and SurroundOcc [51] benchmarks, while using only ResNet50 [12] and a 256x704 input image resolution.
-Specifically, on the Occ3D-nuScenes validation set, DAOcc achieves 53.82 mIoU when using the camera mask during training and 48.2 RayIoU without using the camera mask.
-Additionally, on the SurroundOcc validation set, DAOcc achieves 45.0 IoU and 30.5 mIoU.
+The results show that DAOcc establishes the new state-of-the-art performance on the Occ3D-nuScenes [15] and Occ3D-Waymo [15] benchmarks, while using only ResNet50 [21] with a 256×704 input image resolution.
+Specifically, on the Occ3D-nuScenes validation set, DAOcc achieves a 2.67 mIoU improvement over SGDOCC [7] when trained with camera mask and a 4.7 RayIoU improvement over ALOcc [8] without camera mask.
+Additionally, on the Occ3D-Waymo validation set, DAOcc demonstrates a 6.08 mIoU boost compared to BEVFormer-Fusion [15].
+Furthermore, with TensorRT optimization, DAOcc reaches 104.9 FPS on an NVIDIA RTX 4090 GPU while maintaining 54.2 mIoU on the Occ3D-nuScenes validation set, significantly surpassing the mIoU of previous state-of-the-art methods.
 In conclusion, our contributions are summarized as follows:
 We design a simple yet efficient multi-modal 3D semantic occupancy prediction baseline, eliminating the need for complex deformable attention [39] modules as well as image depth estimation during feature fusion.
-We propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection to assist in achieving superior performance while using a deployment-friendly image encoder and practical input image resolution.
-We introduce a BEV View Range Extension strategy, which provides a larger BEV field of view, offering more contextual information and mitigating the adverse effects of reduced image resolution.
-We establish the new state-of-the-art performance on Occ3D-nuScenes dataset and SurroundOcc dataset, while using only ResNet50 and a 256×704 input image resolution.
-2 Related Work
-3D Occupancy Prediction: 3D occupancy prediction aims to map all occupied voxels in the environment and assign semantic labels, thus providing more fine-grained perception results.
-Camera-based 3D occupancy prediction has gained significant attention due to its cost-effectiveness.
-Early works [4, 16, 62] lack dense surround occupancy supervision, resulting in sparse occupancy predictions.
-To obtain dense occupancy prediction, OpenOccupancy [49], SurroundOcc [51] and Occ3D [44] developed methods for generating dense occupancy labels and established benchmarks on their respective proposed datasets.
-It is worth noting that SurroundOcc does not introduce manual annotation.
-The emergence of these benchmarks further promoted the development of 3D occupancy prediction.
-PanoOcc [50] adopts voxel query and a coarse-to-fine approach to learn a unified occupancy representation.
-FlashOcc [57] uses BEV features with 2D convolution to achieve efficient feature extraction and proposes the Channel-to-Heigh transformation to lift the output of BEV to 3D space.
-SparseOcc [31] leverages a sparse voxel decoder to reconstruct the sparse geometry of a scene and uses sparse queries to predict masks and labels.
-However, all camera-based methods inevitably need to cope with the challenges of adverse lighting and weather conditions, and many works inevitably lack robustness due to the ill-posed monocular depth estimation problem [14, 21].
-While camera-based occupancy prediction has demonstrated promising results, multi-modal approaches offer greater reliability and robustness, making them indispensable for practical applications in autonomous driving and robotics.
+We propose DAOcc, a novel multi-modal occupancy prediction framework that leverages detachable 3D object detection supervision to improve foreground class discrimination and maintain low inference complexity, thus achieving superior performance while using a deployment-friendly image encoder and practical input image resolution.
+We introduce a BEV View Range Extension strategy to mitigate the adverse effects of reduced image resolution, and design a flexible feature processing pipeline consisting of coordinate system transformation and feature interpolation, which eliminates the complex manual feature space alignment.
+Our proposed method, DAOcc, establishes the new state-of-the-art performance on the Occ3D [15] benchmark, while maintaining competitive inference speed.
+With TensorRT optimization, DAOcc can achieve 104.9 FPS while significantly surpassing the mIoU of the previous state-of-the-art methods.
+II Related Work
+II-A 3D Occupancy Prediction
+3D occupancy prediction aims to map all occupied voxels in the environment and assign semantic labels, thus providing more fine-grained perception results.
+While camera-based 3D occupancy prediction [2, 40, 41, 12, 14, 15, 3, 18, 42, 43, 44, 8, 16, 19, 20, 17] has demonstrated promising results, multi-modal approaches offer greater reliability and robustness, making them indispensable for practical applications in autonomous driving and robotics.
 Due to the fact that cameras are susceptible to changing lighting and weather conditions, OccFusion [10] enhances the accuracy and robustness of the occupancy network by integrating features from LiDAR and radar.
@@
 Although OccFusion [11] avoids depth estimation, using deformable attention incurs a greater computational burden [34].
-Hydra [52] extends FB-BEV [25] to the camera-radar fusion domain and improves the accuracy of depth estimation with the help of radar features.
 Co-Occ [13] employs K-nearest neighbor search in selecting neighboring camera features to enhance the corresponding LiDAR features and proposes a regularization based on implicit volume rendering.
 However, this regularization only utilizes the distance ground truth of the point cloud and does not leverage its intrinsic geometry information.
-EFFOcc [43] proposes an efficient and lightweight multi-modal 3D occupancy network, but it significantly relies on the 3D object detection pre-training of the LiDAR branch for optimal results.
-This dependence, however, limits the flexibility of the network structure design.
-Specifically, to maximize the benefits of the 3D detection pre-training, EFFOcc has to adjust its network architecture to align more closely with well-established LiDAR detection networks.
-In this work, we introduce a simple yet efficient multi-modal occupancy prediction network, eliminating the need for complex deformable attention [23] as well as image depth estimation during feature fusion, and does not require 3D object detection pre-training.
-Furthermore, to fully exploit the inherent geometry information in point cloud features, we add supervision of 3D object detection as an auxiliary supervision on the fused features.
-Multi-Modal 3D Object Detection: Recent multi-modal 3D object detection methods [33, 19, 20, 6, 46] mainly focus on learning effective BEV feature representations.
+SDGOCC [7] leverages both the geometric and semantic information of point clouds to guide the 2D-3D view transformation, and introduces a fusion-to-occupancy-driven active distillation module that selectively transfers multimodal knowledge to image features based on LiDAR-identified regions.
+Compared to Co-Occ, SDGOCC additionally introduces point cloud semantic information, but its utilization of point cloud geometry information remains limited—relying solely on sparse point cloud based depth supervision for monocular depth estimation, without addressing the inherent ill-posed problem of this task.
+In this work, we introduce a simple yet efficient multi-modal occupancy prediction network, eliminating the need for complex deformable attention [39] as well as image depth estimation during feature fusion, and avoiding the requirement for 3D object detection pre-training.
+Furthermore, to fully exploit the inherent geometry information in point cloud features, we incorporate 3D object detection as auxiliary supervision on the fused BEV features.
+II-B Multi-Modal 3D Object Detection
+Multi-modal 3D object detection has achieved great success, surpassing camera-based [46, 47] and Lidar-based [48] methods by leveraging semantic and geometric information.
+Recent multi-modal 3D object detection methods [49, 50, 51, 52] mainly focus on learning effective BEV feature representations.
 TransFusion [53] proposes a two-stage transformer-decoder based detection head and applies cross attention to obtain image features for each object query.
@@
 In this work, we incorporate the simple yet effective feature fusion approach of BEVFusion [49] and introduce 3D object detection as an auxiliary branch during training.
-3 Proposed Method
-3.1 Overall Framework
+II-C Detection-Occupancy Interaction
+Recent work has explored tight couplings between 3D occupancy estimation and 3D object detection, with mutual benefits.
+OccupancyM3D [55] uses voxel-wise occupancy as auxiliary dense supervision to enhance monocular 3D detection via richer 3D features and geometry priors.
+SOGDet [56] adds a semantic-occupancy branch that provides dense scene context to multi-view detectors, boosting localization and robustness.
+Zheng et al.
+[57] recently proposed an object-centric occupancy completion approach, focusing on high-resolution modeling of foreground objects.
+DAOcc is complementary to these lines of work, while differing in both supervision direction and practical emphasis.
+III Proposed Method
+III-A Overall Framework
 Our purpose is to fully utilize the inherent geometry and structure information in point cloud features within a multi-modal occupancy prediction framework.
@@
 These modules collectively form the basic network architecture of DAOcc, which will be elaborated in Sec.
-3.2.
+III-B.
 On the foundation of the basic network, we introduce the BVRE strategy (Sec.
-3.3) to compensate for the information loss that arises from the reduction in image resolution.
+III-C) to compensate for the information loss that arises from the reduction in image resolution.
 This strategy aims to enrich the spatial contextual semantic information by broadening the perceptual range from the BEV perspective.
 Furthermore, to fully leverage the inherent geometry and structure information within point cloud features, we incorporate 3D object detection as auxiliary supervision (Sec.
-3.4).
+III-D).
 This auxiliary supervision not only enhances the discriminability of the fused features but also leads to a very concise overall training loss (Sec.
-3.5) for our proposed framework.
-3.2 Basic Network
-LiDAR Encoder.
+III-E) for our proposed framework.
+III-B Basic Network
+III-B1 LiDAR Encoder
 The method of embedding raw LiDAR points into 3D voxelized features is consistent with SECOND [31].
@@
 Next, we apply 3D sparse convolutions [31] to encode these voxel features, generating spatially compressed LiDAR voxel features , where represents the feature dimensions.
-Camera Encoder.
+III-B2 Camera Encoder
 For image feature extraction, taking surround images as input, we first use ResNet50 [21] as the backbone to extract multi-scale features, denoted as , where represent features extracted after downsampling.
@@
 The output feature map has a shape of , where and represent the input resolution of the image, and and indicate the number of channels and the number of surround images, respectively.
-Projection and Interpolation.
+III-B3 Projection and Interpolation
 For image-related occupancy prediction, a key step is to transform image features from 2D image plane to 3D volume space.
-Most existing methods use monocular depth estimation [49, 43, 13, 57] or deformable attention [51, 44, 50, 61].
+Most existing methods use monocular depth estimation [12, 16, 18] or deformable attention [14, 15, 3, 11].
 However, monocular depth estimation is inherently an ill-posed problem [32, 33], while deformable attention imposes a significant computational burden [34].
@@
 Next, the sub-pixel projection positions of the retained points are bilinearly interpolated to generate the image features corresponding to each voxel.
-For voxels located in the overlapping viewing regions of the surround cameras, we average the image features from the two corresponding cameras to obtain the final feature for each voxel.
+For voxels located in the overlapping viewing regions of the surround cameras, we randomly select the image feature from one of the two corresponding cameras to obtain the final feature for each voxel.
 The output camera’s voxel features can be denoted as .
-BEV Encoder.
+III-B4 BEV Encoder
 Given the fused feature , we further refine by passing it through three blocks of ResNet18 [21], resulting in two feature maps, and , extracted at two scales from the first and last blocks respectively.
@@
 The output refined BEV features can be denoted as .
-3.3 BVRE
+III-C BVRE
 As shown in Figure 4, due to the sparsity of point clouds, the increase in computational cost resulting from expanding the processing range of the point cloud is very small compared to increasing the image resolution.
@@
 The function of is to scale the coordinate values to a range from -1 to 1.
-3.4 Detachable Auxiliary Detection Head
+III-D Detachable Auxiliary Detection Head
 To fully exploit the inherent geometry and structure information in point cloud features, we employ a detachable auxiliary detection head to further supervise the feature fusion process, thereby enhancing the information about object boundaries and scene structures in the fused features.
@@
 The total loss for the auxiliary detection head can be denoted as:
-3.5 Overall Objective Function
-Most existing methods rely on complex combinations of loss functions [49, 50, 24, 58, 42, 53, 37, 61, 38, 43], such as a combination of several or all of the focal loss [41], the scene-class affinity loss [4], the dice loss, the lovasz-softmax loss [2], the depth loss [22], etc., to achieve the expected performance.
+III-E Overall Objective Function
+Most existing methods rely on complex combinations of loss functions [12, 3, 17, 62, 19, 20, 10, 11, 13], such as a combination of several or all of the focal loss [36], the scene-class affinity loss [2], the dice loss, the lovasz-softmax loss [37], the depth loss [38], etc., to achieve the expected performance.
 In contrast, our approach utilizes a detachable 3D object detection branch for auxiliary supervision and requires only a simple cross-entropy loss for occupancy prediction.
 The total loss for our framework can be defined as:
-4 Experiments
-Method | mIoU | Modality | Image Size | 2D Backbone |
+IV Experiments
+| Method | mIoU | Modality | Image Size | 2D Backbone | FPS |
 others |
@@
 vegetation |
-BEVDetOcc(2f) [13] | 42.02 | C | 512 1408 | Swin-B | 12.2 | 49.6 | 25.1 | 52.0 | 54.5 | 27.9 | 28.0 | 28.9 | 27.2 | 36.4 | 42.2 | 82.3 | 43.3 | 54.6 | 57.9 | 48.6 | 43.6 |
-PanoOcc(4f) [50] | 42.13 | C | 864 1600 | R101-DCN | 11.7 | 50.5 | 29.6 | 49.4 | 55.5 | 23.3 | 33.3 | 30.6 | 31.0 | 34.4 | 42.6 | 83.3 | 44.2 | 54.4 | 56.0 | 45.9 | 40.4 |
-FB-OCC(16f) [24] | 48.90 | C | 960 1760 | VoVNet-99 | 14.3 | 57.0 | 38.3 | 57.7 | 62.1 | 34.4 | 39.4 | 38.8 | 39.4 | 42.9 | 50.0 | 86.0 | 50.2 | 60.1 | 62.5 | 52.4 | 45.7 |
-FlashOcc(2f) [57] | 43.52 | C | 512 1408 | Swin-B | 13.4 | 51.1 | 27.7 | 51.6 | 56.2 | 27.3 | 30.0 | 29.9 | 29.8 | 37.8 | 43.5 | 83.8 | 46.6 | 56.2 | 59.6 | 50.8 | 44.7 |
-COTR(8f) [36] | 46.21 | C | 512 1408 | Swin-B | 14.9 | 53.3 | 35.2 | 50.8 | 57.3 | 35.4 | 34.1 | 33.5 | 37.1 | 39.0 | 45.0 | 84.5 | 48.7 | 57.6 | 61.1 | 51.6 | 46.7 |
-RadOcc [60] | 49.38 | C+L | 512 1408 | Swin-B | 10.9 | 58.2 | 25.0 | 57.9 | 62.9 | 34.0 | 33.5 | 50.1 | 32.1 | 48.9 | 52.1 | 82.9 | 42.7 | 55.3 | 58.3 | 68.6 | 66.0 |
-OccFusion [37] | 46.67 | C+L+R | 900 1600 | R101-DCN | 12.4 | 50.3 | 31.5 | 57.6 | 58.8 | 34.0 | 41.0 | 47.2 | 29.7 | 42.0 | 48.0 | 78.4 | 35.7 | 47.3 | 52.7 | 63.5 | 63.3 |
-OccFusion [61] | 48.74 | C+L | 900 1600 | R101 | 12.4 | 51.8 | 33.0 | 54.6 | 57.7 | 34.0 | 43.0 | 48.4 | 35.5 | 41.2 | 48.6 | 83.0 | 44.7 | 57.1 | 60.0 | 62.5 | 61.3 |
-HyDRa [52] | 44.4 | C+R | 256 704 | R50 | - | - | - | 52.3 | 56.3 | - | 35.9 | 35.1 | - | - | 44.1 | - | - | - | - | - | - |
-EFFOcc [43] | 52.62 | C+L | 512 1408 | Swin-B | 14.7 | 58.6 | 34.7 | 60.5 | 65.5 | 36.0 | 40.4 | 51.5 | 41.0 | 48.5 | 54.6 | 84.2 | 46.4 | 57.9 | 60.8 | 70.9 | 68.4 |
-DAOcc (Ours) | 53.82 | C+L | 256 704 | R50 | 12.4 | 59.6 | 38.4 | 61.9 | 67.1 | 35.3 | 48.2 | 59.1 | 43.5 | 50.9 | 56.3 | 83.0 | 44.7 | 56.7 | 59.9 | 70.0 | 68.1 |
-Method | RayIoU | Modality | Image Size | 2D Backbone | RayIoU1m | RayIoU2m | RayIoU4m | mIoU |
-BEVFormer(4f) [23] | 32.4 | C | 900 1600 | R101 | 26.1 | 32.9 | 38.0 | 39.2 |
-RenderOcc [39] | 19.5 | C | 512 1408 | Swin-B | 13.4 | 19.6 | 25.5 | 24.4 |
-SimpleOcc [9] | 22.5 | C | 336 672 | R101 | 17.0 | 22.7 | 27.9 | 31.8 |
-BEVDetOcc(8f) [13] | 32.6 | C | 384 704 | R50 | 26.6 | 33.1 | 38.2 | 39.3 |
-FB-Occ(16f) [24] | 33.5 | C | 256 704 | R50 | 26.7 | 34.1 | 39.7 | 39.1 |
-SparseOcc(16f) [31] | 36.1 | C | 256 704 | R50 | 30.2 | 36.8 | 41.2 | 30.9 |
-Panoptic-FlashOcc [58] | 35.2 | C | 256 704 | R50 | 29.4 | 36.0 | 40.1 | 29.4 |
-Panoptic-FlashOcc(8f) [58] | 38.5 | C | 256 704 | R50 | 32.8 | 39.3 | 43.4 | 31.6 |
-DAOcc (Ours) | 48.2 | C+L | 256 704 | R50 | 44.4 | 48.7 | 51.6 | 47.9 |
-Method | IoU | mIoU | Modality | Image Size | 2D Backbone |
+| BEVDetOcc(2F) [16] | 42.02 | C | 512 1408 | Swin-B | 1.2 | 12.2 | 49.6 | 25.1 | 52.0 | 54.5 | 27.9 | 28.0 | 28.9 | 27.2 | 36.4 | 42.2 | 82.3 | 43.3 | 54.6 | 57.9 | 48.6 | 43.6 |
+| PanoOcc(4F) [3] | 42.13 | C | 864 1600 | R101-DCN | 3.0 | 11.7 | 50.5 | 29.6 | 49.4 | 55.5 | 23.3 | 33.3 | 30.6 | 31.0 | 34.4 | 42.6 | 83.3 | 44.2 | 54.4 | 56.0 | 45.9 | 40.4 |
+| FB-OCC(16F) [17] | 48.90 | C | 960 1760 | VoVNet-99 | - | 14.3 | 57.0 | 38.3 | 57.7 | 62.1 | 34.4 | 39.4 | 38.8 | 39.4 | 42.9 | 50.0 | 86.0 | 50.2 | 60.1 | 62.5 | 52.4 | 45.7 |
+| FlashOcc(2F)* [18] | 43.52 | C | 512 1408 | Swin-B | 1.2 | 13.4 | 51.1 | 27.7 | 51.6 | 56.2 | 27.3 | 30.0 | 29.9 | 29.8 | 37.8 | 43.5 | 83.8 | 46.6 | 56.2 | 59.6 | 50.8 | 44.7 |
+| COTR(8F) [43] | 46.21 | C | 512 1408 | Swin-B | 0.3 | 14.9 | 53.3 | 35.2 | 50.8 | 57.3 | 35.4 | 34.1 | 33.5 | 37.1 | 39.0 | 45.0 | 84.5 | 48.7 | 57.6 | 61.1 | 51.6 | 46.7 |
+| STCOcc(16F) [44] | 45.0 | C | 512 1408 | R50 | 1.4 | 15.2 | 52.3 | 32.2 | 50.5 | 56.5 | 31.7 | 33.9 | 33.4 | 33.8 | 38.9 | 44.9 | 83.9 | 47.4 | 57.1 | 60.1 | 50.6 | 42.7 |
+| ALOcc(16F) [8] | 50.6 | C | 512 1408 | Swin-B | 1.5 | 17.0 | 58.3 | 39.7 | 56.6 | 63.2 | 33.2 | 41.3 | 40.3 | 40.8 | 43.8 | 51.0 | 87.0 | 52.7 | 62.0 | 65.2 | 57.7 | 50.9 |
+| RadOcc [9] | 49.38 | C+L | 512 1408 | Swin-B | - | 10.9 | 58.2 | 25.0 | 57.9 | 62.9 | 34.0 | 33.5 | 50.1 | 32.1 | 48.9 | 52.1 | 82.9 | 42.7 | 55.3 | 58.3 | 68.6 | 66.0 |
+| OccFusion [10] | 46.67 | C+L+R | 900 1600 | R101-DCN | 1.0 | 12.4 | 50.3 | 31.5 | 57.6 | 58.8 | 34.0 | 41.0 | 47.2 | 29.7 | 42.0 | 48.0 | 78.4 | 35.7 | 47.3 | 52.7 | 63.5 | 63.3 |
+| OccFusion [11] | 48.74 | C+L | 900 1600 | R101 | - | 12.4 | 51.8 | 33.0 | 54.6 | 57.7 | 34.0 | 43.0 | 48.4 | 35.5 | 41.2 | 48.6 | 83.0 | 44.7 | 57.1 | 60.0 | 62.5 | 61.3 |
+| SDGOCC [7] | 51.66 | C+L | - | R50 | 7.5 | 13.2 | 57.8 | 24.3 | 60.3 | 64.3 | 36.2 | 39.4 | 52.4 | 35.8 | 50.9 | 53.7 | 84.6 | 47.5 | 58.0 | 61.6 | 70.7 | 67.7 |
+| DAOcc (Ours) | 54.33 | C+L | 256 704 | R50 | 7.8 | 13.0 | 60.7 | 39.8 | 64.0 | 66.5 | 36.3 | 49.0 | 60.1 | 44.3 | 50.7 | 55.9 | 82.9 | 44.6 | 56.8 | 60.6 | 70.1 | 68.3 |
+| Method | Framework | Precision | Hardware | mIoU | FPS |
+| DAOcc | PyTorch | FP32 | RTX4090 | 54.33 | 7.8 |
+| DAOcc | TensorRT | FP16 | RTX4090 | 54.20 | 104.9 |
+| DAOcc | TensorRT | FP16+INT8 | Orin | 53.70 | 20.0 |
+| Method | mIoU | RayIoU | Modality | Image Size | 2D Backbone | FPS | RayIoU1m | RayIoU2m | RayIoU4m | |
+| train | val | |||||||||
+| BEVFormer(4F) [39] | 39.2 | 32.4 | - | C | 900 1600 | R101 | - | 26.1 | 32.9 | 38.0 |
+| RenderOcc [63] | 24.4 | 19.5 | - | C | 512 1408 | Swin-B | - | 13.4 | 19.6 | 25.5 |
+| SimpleOcc [64] | 31.8 | 22.5 | - | C | 336 672 | R101 | - | 17.0 | 22.7 | 27.9 |
+| BEVDetOcc(8F) [16] | 39.3 | 32.6 | C+L | C | 384 704 | R50 | - | 26.6 | 33.1 | 38.2 |
+| FB-Occ(16F) [17] | 39.1 | 33.5 | C+L | C | 256 704 | R50 | - | 26.7 | 34.1 | 39.7 |
+| SparseOcc(16F) [42] | 30.9 | 36.1 | - | C | 256 704 | R50 | - | 30.2 | 36.8 | 41.2 |
+| Panoptic-FlashOcc [62] | 29.4 | 35.2 | C+L | C | 256 704 | R50 | - | 29.4 | 36.0 | 40.1 |
+| Panoptic-FlashOcc(8F) [62] | 31.6 | 38.5 | C+L | C | 256 704 | R50 | - | 32.8 | 39.3 | 43.4 |
+| STCOcc(16F) [44] | - | 42.1 | C+L | C | 512 1408 | R50 | 1.4 | 36.9 | 42.8 | 46.7 |
+| ALOcc(16F) [8] | 38.0 | 43.7 | C+L | C | 256 704 | R50 | 6.0 | 37.8 | 44.7 | 48.8 |
+| DAOcc (Ours) | 48.3 | 48.4 | C+L | C+L | 256 704 | R50 | 7.8 | 44.5 | 48.9 | 51.9 |
+| BEV View Range (m) | BEV Voxel Size (m) | H-Range (m) | mIoU | |
+| (a) | 0.075 0.075 | [-6, 4] | 50.57 | |
+| (b) | 0.075 0.075 | [-4, 2] | 50.27 | |
+| (c) | 0.075 0.075 | [-5, 3] | 50.76 | |
+| (d) | 0.075 0.075 | [-5, 3] | 50.98 | |
+| (e) | 0.1 0.1 | [-5, 3] | 49.75 | |
+| (f) | 0.1 0.1 | [-5, 3] | 49.98 | |
+| (g) | 0.075 0.075 | [-5, 3] | 51.26 |
+ADH | mIoU |
+others |
 barrier |
@@
 vegetation |
-SurroundOcc [51] | 31.4 | 20.3 | C | 900 1600 | R101-DCN | 20.5 | 11.6 | 28.1 | 30.8 | 10.7 | 15.1 | 14.0 | 12.0 | 14.3 | 22.2 | 37.2 | 23.7 | 24.4 | 22.7 | 14.8 | 21.8 |
-OccFormer [62] | 29.9 | 20.1 | C | 896 1600 | R101 | 21.1 | 11.3 | 28.2 | 30.3 | 10.6 | 15.7 | 14.4 | 11.2 | 14.0 | 22.6 | 37.3 | 22.4 | 24.9 | 23.5 | 15.2 | 21.1 |
-C-CONet [49] | 26.1 | 18.4 | C | 896 1600 | R101 | 18.6 | 10.0 | 26.4 | 27.4 | 8.6 | 15.7 | 13.3 | 9.7 | 10.9 | 20.2 | 33.0 | 20.7 | 21.4 | 21.8 | 14.7 | 21.3 |
-FB-Occ [24] | 31.5 | 19.6 | C | 896 1600 | R101 | 20.6 | 11.3 | 26.9 | 29.8 | 10.4 | 13.6 | 13.7 | 11.4 | 11.5 | 20.6 | 38.2 | 21.5 | 24.6 | 22.7 | 14.8 | 21.6 |
-RenderOcc [39] | 29.2 | 19.0 | C | 896 1600 | R101 | 19.7 | 11.2 | 28.1 | 28.2 | 9.8 | 14.7 | 11.8 | 11.9 | 13.1 | 20.1 | 33.2 | 21.3 | 22.6 | 22.3 | 15.3 | 20.9 |
-LMSCNet [40] | 36.6 | 14.9 | L | - | - | 13.1 | 4.5 | 14.7 | 22.1 | 12.6 | 4.2 | 7.2 | 7.1 | 12.2 | 11.5 | 26.3 | 14.3 | 21.1 | 15.2 | 18.5 | 34.2 |
-L-CONet [49] | 39.4 | 17.7 | L | - | - | 19.2 | 4.0 | 15.1 | 26.9 | 6.2 | 3.8 | 6.8 | 6.0 | 14.1 | 13.1 | 39.7 | 19.1 | 24.0 | 23.9 | 25.1 | 35.7 |
-M-CONet [49] | 39.2 | 24.7 | C+L | 896 1600 | R101 | 24.8 | 13.0 | 31.6 | 34.8 | 14.6 | 18.0 | 20.0 | 14.7 | 20.0 | 26.6 | 39.2 | 22.8 | 26.1 | 26.0 | 26.0 | 37.1 |
-Co-Occ [38] | 41.1 | 27.1 | C+L | 896 1600 | R101 | 28.1 | 16.1 | 34.0 | 37.2 | 17.0 | 21.6 | 20.8 | 15.9 | 21.9 | 28.7 | 42.3 | 25.4 | 29.1 | 28.6 | 28.2 | 38.0 |
-OccFusion [37] | 44.7 | 27.3 | C+L+R | 900 1600 | R101-DCN | 27.1 | 19.6 | 33.7 | 36.2 | 21.7 | 24.8 | 25.3 | 16.3 | 21.8 | 30.0 | 39.5 | 19.9 | 24.9 | 26.5 | 28.9 | 40.4 |
-DAOcc (Ours) | 45.0 | 30.5 | C+L | 256 704 | R50 | 30.8 | 19.5 | 34.0 | 38.8 | 25.0 | 27.7 | 29.9 | 22.5 | 23.2 | 31.6 | 41.0 | 25.9 | 29.4 | 29.9 | 35.2 | 43.5 |
-4.1 Datasets and Evaluation Metrics
+| ✓ | 51.26 | 13.01 | 56.94 | 20.6 | 61.49 | 65.51 | 35.08 | 38.89 | 55.44 | 36.65 | 48.46 | 53.82 | 83.22 | 45.08 | 57.35 | 60.63 | 70.83 | 68.47 | |||||||||||||||||||||||||||||||||||||
+| Scheduler | mIoU |
+| EP15 | 52.82 |
+| EP24 | 53.39 |
+| CBGS(EP6) | 53.82 |
+| CBGS(EP6)+LR2 | 54.33 |
+| Method | mIoU | Modality | Image Size | 2D Backbone |
+GO |
+vehicle |
+bicyclist |
+pedestrian |
+sign |
+traffic light |
+pole |
+Cons. Cone |
+bicycle |
+motorcycle |
+building |
+vegetation |
+tree trunk |
+road |
+sidewalk |
+| BEVDet [16] | 9.88 | C | - | - | 0.13 | 13.06 | 2.17 | 10.15 | 7.80 | 5.85 | 4.62 | 0.94 | 1.49 | 0.0 | 7.27 | 10.06 | 2.35 | 48.15 | 34.12 |
+| TPVFormer [40] | 16.76 | C | - | - | 3.89 | 17.86 | 12.03 | 5.67 | 13.64 | 8.49 | 8.90 | 9.95 | 14.79 | 0.32 | 13.82 | 11.44 | 5.8 | 73.3 | 51.49 |
+| BEVFormer [39] | 16.76 | C | - | - | 3.48 | 17.18 | 13.87 | 5.9 | 13.84 | 2.7 | 9.82 | 12.2 | 13.99 | 0.0 | 13.38 | 11.66 | 6.73 | 74.97 | 51.61 |
+| CTF-Occ [15] | 18.73 | C | 640 960 | R101-DCN | 6.26 | 28.09 | 14.66 | 8.22 | 15.44 | 10.53 | 11.78 | 13.62 | 16.45 | 0.65 | 18.63 | 17.3 | 8.29 | 67.99 | 42.98 |
+| OPUS [65] | 19.00 | C | - | - | 4.66 | 27.07 | 19.39 | 6.53 | 18.66 | 6.41 | 11.44 | 10.40 | 12.90 | 0.00 | 18.73 | 18.11 | 7.46 | 72.86 | 50.31 |
+| ODG [66] | 21.35 | C | 640 960 | R50 | 5.09 | 31.34 | 22.4 | 19.06 | 15.24 | 6.09 | 12.51 | 12.77 | 13.59 | 0.00 | 21.49 | 17.89 | 8.37 | 78.19 | 56.28 |
+| LiDAR-Only [15] | 29.74 | L | - | - | 1.01 | 57.41 | 35.31 | 20.33 | 11.7 | 13.01 | 36.21 | 7.81 | 0.13 | 0.0 | 57.83 | 54.71 | 27.07 | 69.15 | 54.47 |
+| BEVFormer-Fusion [15] | 39.05 | C+L | - | - | 5.11 | 64.61 | 52.35 | 21.52 | 32.74 | 17.1 | 42.62 | 27.75 | 13.36 | 0.05 | 63.65 | 60.51 | 35.64 | 81.89 | 66.84 |
+| DAOcc (Ours) | 44.69 | C+L | 256 704 | R50 | 7.64 | 65.60 | 54.05 | 43.71 | 52.78 | 32.05 | 46.96 | 32.68 | 31.68 | 0.68 | 55.94 | 60.19 | 42.59 | 79.62 | 64.09 |
+| DAOcc* (Ours) | 45.13 | C+L | 256 704 | R50 | 8.15 | 65.90 | 54.40 | 43.36 | 52.86 | 32.93 | 47.54 | 33.54 | 32.29 | 0.67 | 57.49 | 60.56 | 42.86 | 79.78 | 64.69 |
+IV-A Datasets and Evaluation Metrics
+IV-A1 Occ3D-nuScenes
 Occ3D-nuScenes [15] is a large-scale benchmark dataset for 3D occupancy prediction, which contains 700 scenes for training and 150 scenes for validation.
@@
 This range is discretized into voxels, each with a size of [0.4m, 0.4m, 0.4m].
-Each occupied voxel is assigned one of 17 semantic labels, including 16 common classes and a general object class labeled as ‘others’.
+Each occupied voxel is assigned one of 17 semantic classes, including 16 common classes and a general object class labeled as ‘others’.
 The dataset includes RGB images captured by six surrounding cameras and LiDAR point clouds, enabling dense voxel-wise annotations for 3D scene understanding.
 Additionally, the dataset provides visibility masks for both LiDAR and camera modes, which can be utilized for training purposes.
-SurroundOcc [51] is an automatically generated occupancy prediction dataset that does not require extra human annotation.
-It is built upon the nuScenes [3] and generates dense occupancy ground truth by utilizing existing 3D detection and 3D semantic segmentation labels.
-The dataset uses the nuScenes training set to train the model and employs the nuScenes validation set for evaluation purposes.
-The occupancy prediction range is set to [-50m, 50m] for the and axes, and [-5m, 3m] for the -axis, all within the LiDAR coordinate system.
-The final output 3D occupancy grid has a shape of 200×200×16, with each voxel measuring [0.5m, 0.5m, 0.5m].
-Evaluation metrics: Following previous methods, we use the mean Intersection over Union (mIoU) as our evaluation metric when training models with camera visible mask provided by Occ3D-nuScenes, ensuring a fair comparison with existing approaches.
-When the camera visible mask is excluded from the training phase, we report an additional semantic segmentation metric, RayIoU [31], which has been recently introduced.
-For the SurroundOcc dataset, we report semantic segmentation performance using IoU and mIoU.
-Here, IoU is defined as the intersection over union of occupied voxels, calculated while ignoring their semantic categories, and it is used as an evaluation metric for the scene completion (SC) task.
-4.2 Implementation Details
+IV-A2 Occ3D-Waymo
+Occ3D-Waymo [15] is a large-scale benchmark dataset for 3D occupancy prediction, containing 798 training scenes and 202 validation scenes.
+It shares the same perception range and voxel size as Occ3D-nuScenes.
+Each occupied voxel is annotated with one of the 15 semantic classes, including a General Object (GO) class.
+The dataset provides RGB images from five cameras (with raw resolutions of for front, front-left, and front-right cameras, and for side-left and side-right cameras) and LiDAR point clouds, along with visibility masks for LiDAR and camera modalities to facilitate training.
+IV-A3 Evaluation metrics
+Following previous methods, we use the mean Intersection over Union (mIoU) as our evaluation metric when training models with the camera visible mask provided by Occ3D-nuScenes and Occ3D-Waymo, ensuring a fair comparison with existing approaches.
+When the camera visible mask is excluded from the training phase, we report an additional semantic segmentation metric, RayIoU [42], which is a recently introduced metric.
+RayIoU evaluates occupancy along query rays at three distance thresholds (1m, 2m, 4m), with the mean of these results serving as the final score.
+IV-B Implementation Details
 We use ResNet-50 [21], pre-trained on nuImages [24] by MMDetection [67], as our image backbone and adjust the input image resolution to 256×704.
@@
 Unless conducting ablation experiments, we train our DAOcc using CBGS [69] for 6 epochs, the input point cloud range is set to [-54.0, -54.0, -5.0, 54.0, 54.0, 3.0] and the voxel size is set to [0.075, 0.075, 0.2].
-All models are trained on 8 NVIDIA RTX4090 GPUs with a batch size of 4.
+All models are trained on 4 NVIDIA RTX4090 GPUs with a batch size of 4.
 Additionally, if the camera visible mask is not used during training, we randomly discard 80% of empty voxels, consistent with Zhao et al.
 [70].
-Temporal information fusion, exponential moving average (EMA) for weight parameter and test-time augmentation is not used.
-4.3 Comparison with State-of-the-Art Methods
-Method | Modality | 3D Backbone | Param | mIoU |
-FB-Occ [24] | C | - | 130.8M | 48.90 |
-RadOcc [60] | C+L | SECOND | 135.4M | 49.38 |
-EFFOcc [43] | C+L | SECOND | 108.6M | 52.62 |
-DAOcc | C+L | SECOND | 56.2M | 53.82 |
-Comparison on Occ3D-nuScenes with camera mask.
-As presented in Table 1, we compare our proposed DAOcc with both image-based and multi-modal methods on the Occ3D-nuScenes dataset.
-All results are provided either directly by their authors or based on their official code implementations.
-Our DAOcc establishes a new state-of-the-art performance, achieving an outstanding 53.82 mIoU, while utilizing a 2D backbone ResNet50 [12] and 256×704 input image resolution.
-Compared with the latest state-of-the-art multi-modal method EFFOcc [43], DAOcc achieves improvement of 1.2 mIoU.
-Notably, the image feature extraction network Swin-B [32] in EFFOcc has approximately 3 times the number of parameters of ResNet50, and EFFOcc uses a larger input image resolution of 512×1408.
-This highlights our method’s ability to mine more useful information from point cloud data.
-In addition, to achieve optimal results, EFFOcc requires the LiDAR branch to undergo 3D object detection pre-training, which leads to an improvement of approximately 3 mIoU.
-In contrast, DAOcc does not require LiDAR branch pre-training, but instead seamlessly integrates 3D object detection as an auxiliary branch into the multi-modal occupancy prediction framework, simplifying the model training process and achieving better performance.
-Furthermore, we compare the complexity of the models, as shown in Table 4.
-Compared to the two top-performing multi-modal methods, RadOcc [60] and EFFOcc [43], as well as the top-performing image-based method FB-Occ [24], our proposed DAOcc has only about half the number of parameters but achieves significantly better performance (outperforming RadOcc, EFFOcc, and FB-Occ by 4.44, 1.2 and 4.92 mIoU).
-Comparison on Occ3D-nuScenes without camera mask.
-To the best of our knowledge, no existing multi-modal occupancy prediction work has used RayIoU metrics to report model performance, so we mainly compare with image-based methods.
-As shown in Table 2, benefiting from the effective use of multi-modal input, especially point cloud data, our DAOcc achieves an outstanding 48.2 RayIoU and significantly outperforms image-based methods by a large margin, even though many of these methods use multi-frame temporal fusion.
-It should be noted that some image-based methods, such as Panoptic-FlashOcc and FB-Occ [24], adopt the Explicit Depth Supervision [22] and leverage depth information derived from point cloud data to supervise the training of monocular depth estimation from images, so these methods can also be regarded as multi-modal during the training phase.
-Comparison on SurroundOcc.
-Table 3 presents a quantitative comparison on the SurroundOcc validation set, showcasing the performance of our DAOcc against other methods.
-Leveraging both camera and LiDAR inputs, DAOcc achieves a notable 3.2 mIoU improvement over OccFusion [37], which incorporates radar data in addition to camera and LiDAR inputs.
-It is worth noting that OccFusion utilizes R101-DCN with a high-resolution input of 900×1600, whereas our method employs a more lightweight Resnet50 backbone and a lower input resolution of 256×704.
-This underscores the effectiveness and efficiency of our proposed method.
-4.4 Ablation Studies
-Ablation studies are conducted on the validation set of Occ3D-nuScenes, with the camera visible mask being used during training.
+Temporal information fusion and test-time augmentation are not used in any experiment.
+Unless otherwise stated, exponential moving average (EMA) for weight parameter is not used.
+IV-C Comparison with State-of-the-Art Methods
+IV-C1 Comparison on Occ3D-nuScenes with camera mask
+As presented in Table I, we compare our proposed DAOcc with both image-based and multi-modal methods on the Occ3D-nuScenes dataset.
+All results are either provided directly by their authors or obtained from their official code implementations.
+FPS for SDGOCC and ALOcc are sourced directly from their original publications, evaluated on NVIDIA RTX 4090.
+All other FPS results are obtained through our own benchmarking on identical hardware (NVIDIA RTX 4090) using the PyTorch FP32 backend.
+Our proposed DAOcc establishes a new state-of-the-art performance, achieving an impressive 54.33 mIoU with only a 2D ResNet50 [21] backbone and a low input image resolution of 256×704, while maintaining a competitive inference speed of 7.8 FPS.
+Compared with the latest state-of-the-art multi-modal method SDGOCC, DAOcc achieves an improvement of 2.67 mIoU and maintains a higher FPS (7.8 vs 7.5).
+Furthermore, DAOcc performs strongly across nearly all foreground semantic categories, which are critical for ensuring autonomous driving safety.
+As shown in Table II, we further evaluate the performance and inference speed of DAOcc under different frameworks and hardware configurations.
+Unlike most LSS-based methods that rely on custom CUDA operators (e.g., BEV Pool [49]), the image branch of DAOcc does not require any custom operators when deployed with TensorRT and can be directly exported to the ONNX format, greatly simplifying the deployment process.
+With TensorRT optimization, DAOcc achieves 104.9 FPS on an RTX 4090 while maintaining 54.2 mIoU, and 20.0 FPS on a Jetson AGX Orin (64GB) while maintaining 53.7 mIoU.
+Although quantization introduces a slight drop in mIoU performance, the resulting 53.7 mIoU still significantly surpasses all other methods reported in Table I.
+IV-C2 Comparison on Occ3D-nuScenes without camera mask
+To the best of our knowledge, no existing multi-modal occupancy prediction methods have reported their performance using the RayIoU metric, so we mainly compare with image-based methods.
+It is worth noting that some image-based methods [16, 17, 62, 44, 8], such as the current state-of-the-art STCOcc and ALOcc, are based on Explicit Depth Supervision [38] and leverage depth information derived from point clouds to supervise monocular depth estimation during training.
+As such, these methods can also be considered multi-modal in the training phase.
+As shown in Table III, benefiting from the effective use of multi-modal inputs, especially point cloud data, our DAOcc achieves 48.3 mIoU and 48.4 RayIoU, significantly outperforming image-based methods by a large margin (surpassing ALOcc by 10.3 mIoU and 4.7 RayIoU), despite many of these methods employing multi-frame temporal fusion, which is an effective technique to improve RayIoU scores (e.g., +3.3 RayIoU for Panoptic-FlashOcc with 8-frame fusion vs.
+single-frame).
+Furthermore, when using the same 2D backbone (ResNet50) and image input resolution (256×704), although DAOcc additionally inputs point cloud data, its inference speed is still faster than the image-based method ALOcc (7.8 vs.
+6.0).
+IV-D Ablation Studies
+Ablation studies are conducted on the validation set of Occ3D-nuScenes and Occ3D-Waymo, with the camera visible mask being used during training.
 Unless otherwise specified, all experiments presented in this section are trained for 15 epochs without employing CBGS [69].
-BEV View Range (m) | BEV Voxel Size (m) | mIoU | |
-(a) | 0.075 0.075 | 50.76 | |
-(b) | 0.075 0.075 | 50.98 | |
-(c) | 0.1 0.1 | 49.75 | |
-(d) | 0.1 0.1 | 49.98 | |
-(e) | 0.075 0.075 | 51.26 |
-Effectiveness of the BVRE.
+IV-D1 Effectiveness of the BVRE
 We propose the BVRE strategy to enrich spatial contextual information by expanding the perceptual range from a BEV (Bird’s Eye View) perspective.
-As shown in Table 5, a comparison of (a) and (b) indicates that, with the same voxel size, expanding the BEV range from [-41.4m, 41.4m] to [-45.6m, 45.6m] yields a performance improvement of 0.22 mIoU.
-Furthermore, the comparison of (a) and (e) shows that further expanding the BEV range to [-54.0m, 54.0m] augments the performance gain to 0.5 mIoU.
+As shown in Table IV, a comparison of (c) and (d) indicates that, with the same voxel size of and the H-Range of [-5m, 3m], expanding the BEV range from [-41.4m, 41.4m] to [-45.6m, 45.6m] yields a performance improvement of 0.22 mIoU (from 50.76 to 50.98).
+Furthermore, the comparison of (c) and (g) shows that further expanding the BEV range to [-54.0m, 54.0m] with the same voxel size and H-Range increases the performance gain to 0.5 mIoU (from 50.76 to 51.26).
 This observation validates the value of spatial context information in a larger BEV range.
-In addition, by comparing (b) and (c), we can find that the size of the voxel has a great impact on the model’s performance. To achieve better performance, we adopt the setting (e) in DAOcc.
-camera mask | ADH | mIoU |
-bicycle |
-motorcycle |
-pedestrian |
-traffic cone |
-✓ | 51.26 | 20.60 | 38.89 | 55.44 | 36.65 | |
-✓ | ✓ | 52.82 (+1.56) | 36.35 (+15.75) | 45.48 (+6.59) | 58.27 (+2.83) | 41.98 (+5.33) |
-44.31 | 14.94 | 32.22 | 43.77 | 34.07 | ||
-✓ | 46.89 (+2.58) | 32.51 (+17.57) | 42.89 (+10.67) | 53.41 (+9.64) | 40.55 (+6.48) |
-Ablation on the Auxiliary Detection Head.
-To fully leverage the inherent geometry and structure information within point cloud features, we incorporate 3D object detection as auxiliary supervision.
-As shown in Table 6, by adding auxiliary 3D object detection supervision, we can achieve an improvement of 1.56 mIoU when using camera visible masks during training and 2.58 mIoU when not using the mask, respectively.
-This demonstrates that adding a detection branch can significantly enhance the discriminability of the multi-modal features.
-Furthermore, we report on the four categories that benefit the most from the 3D object detection supervision.
-Notably, all four categories are foreground objects, which further validates the effectiveness and rationale behind the performance improvement achieved by our method.
-Specifically, when the camera visible mask is used during training, the addition of the auxiliary detection branch results in improvements of 15.75 mIoU, 6.59 mIoU, 2.83 mIoU, and 5.33 mIoU for bicycles, motorcycles, pedestrians, and traffic cones, respectively.
-Conversely, when the camera visible mask is not used during training, the improvements are 17.57 mIoU, 10.67 mIoU, 9.64 mIoU, and 6.48 mIoU for the same categories, respectively.
-5 Conclusion
-In this paper, we propose a novel multi-modal occupancy prediction framework, DAOcc, which aims to achieve superior performance while utilizing a deployment-friendly image backbone and practical input image resolutions.
-DAOcc achieves this superior performance by fully leveraging the unique advantages of point cloud data within a multi-modal framework, advantages that have been largely overlooked by previous methods.
-Extensive experiments conducted on the Occ3D-nuScenes and SurroundOcc datasets demonstrate the superiority of DAOcc compared to existing methods.
-We believe that DAOcc has practical application potential and could inspire future research in multi-modal occupancy prediction.
-Supplementary Material
-Method | Modality | GPU Type | FPS | mIoU |
-FlashOcc(2f) [57] | C | A100-PCIE-40G | 0.9 | 43.52 |
-RadOcc [60] | C+L | - | 0.3 | 49.38 |
-EFFOcc [43] | C+L | RTX A6000 | 1.8 | 52.62 |
-DAOcc(ours) | C+L | RTX 3070Ti | 2.7 | 53.82 |
-6 Runtime Comparisons
-We compare the runtime of different methods, as shown in Table 7.
-The FPS is obtained by testing on the corresponding GPU Type listed in the table.
-Since the source code for RadOcc is unavailable and EFFOcc has not yet released its implementation, we rely on the runtime speeds reported in their respective papers as benchmarks.
-It’s worth noting that the computing capability of the RTX 3070 Ti is significantly lower than that of both the RTX A6000 and the A100.
-The results demonstrate that our proposed DAOcc achieves the best performance while maintaining the fastest runtime speed.
-7 Ablation for Achieving Final Performance
-In addition to the ablation studies presented in Section 4.4, this section provides further ablation experiments.
-Table 8 presents the ablation experiments that contributed to the final performance.
-First, by extending the training time from 15 epochs to 24 epochs, we can observe an improvement of 0.57 mIoU.
-Next, the use of CBGS [65] yields an additional 0.43 mIoU improvement.
-Since CBGS involves resampling, which increases the number of samples per epoch, we limit the training to 6 epochs to maintain a comparable total training time to the 24 epochs without CBGS for a fair comparison.
-8 Detection Performance
-Since our proposed DAOcc incorporates an auxiliary 3D object detection branch into the occupancy prediction framework, we also evaluate the performance of this auxiliary branch.
-The auxiliary detection branch is designed to identify 10 object categories, including cars, trucks, buses, trailers, construction vehicles, pedestrians, motorcycles, bicycles, traffic cones, and barriers.
-All experiments are conducted using the nuScenes [3] dataset for the auxiliary 3D object detection training, and the results are evaluated using the mean average precision (mAP) and nuScenes detection score (NDS) metrics provided by the nuScenes benchmark.
-As shown in Table 9, when DAOcc is trained on the Occ3D-nuScenes [44] dataset (for occupancy prediction) and the camera visible mask is used during training, the auxiliary detection branch achieves a performance of 59.4 mAP and 64.3 NDS.
-Notably, comparable performance is observed when the camera visible mask is not used.
-However, when the SurroundOcc [51] dataset is employed for occupancy prediction training, the auxiliary detection branch achieves a lower performance of 53.6 mAP and 60.2 NDS.
-We attribute this performance gap to the fact that the surroundocc dataset is automatically generated without manual calibration, potentially introducing more noise and negatively affecting model convergence.
-EP15 | EP24 | CBGS (EP6) | mIoU |
-✓ | 52.82 | ||
-✓ | 53.39 | ||
-✓ | 53.82 |
-9 Visualization
+In addition, by comparing (d) and (e), the voxel size has a great impact on the model’s performance.
+Specifically, increasing the voxel size from to while keeping the BEV range and H-Range the same leads to a decrease in mIoU from 50.98 to 49.75.
+Moreover, by analyzing the data in (a), (b), and (c), the H-Range of [-5m, 3m] is a relatively optimal setting, as it consistently achieves higher mIoU values compared to other H-Ranges.
+To achieve better performance, we adopt the setting (g) in DAOcc, which combines the largest BEV range of [-54.0m, 54.0m], the voxel size of m, and the H-Range of [-5m, 3m].
+IV-D2 Ablation on the Auxiliary Detection Head
+To fully leverage the inherent geometry and structure information in point cloud features, we incorporate 3D object detection as auxiliary supervision.
+As shown in Table V, adding auxiliary 3D detection supervision leads to a 1.56 mIoU improvement when training with camera visible masks.
+Notably, this improvement originates entirely from foreground categories, which are highlighted in light gray in Table V.
+These categories correspond to the ones supervised by the auxiliary detection branch, demonstrating the effectiveness of our proposed supervision strategy.
+Meanwhile, we observe that categories without detection supervision all exhibit varying degrees of performance degradation.
+We infer that this is because detection supervision emphasizes the localization of foreground objects, which strengthens the representation of foreground objects and thus weakens that of background categories.
+To mitigate this issue, we believe that additional BEV segmentation supervision could be introduced to balance the representation of foreground and background categories.
+To further investigate how auxiliary 3D detection supervision affects the BEV feature distribution, we visualize intermediate features in Figure 5.
+Subfigures (a) and (b) show the input point cloud and the corresponding ground-truth occupancy map, while (c) and (d) present the BEV features without and with auxiliary 3D detection supervision, respectively.
+The comparison reveals that incorporating auxiliary 3D detection enhances the preservation of structural scene details and significantly improves the distinguishability of foreground objects.
+IV-D3 Ablation on Loss Weight and LiDAR Sweeps
+Figure 6 illustrates the impacts of two factors: the loss weight and LiDAR sweeps on the mIoU.
+To analyze the effect of loss weight , which balances the loss contributions between 3D occupancy prediction and 3D object detection, we vary its value across 0.1, 0.01, 0.001, and 0.0001.
+As shown in the left plot, mIoU rises sharply from 48.17 to 52.82 as decreases from 0.1 to 0.01, then gradually declines to 50.81 as is further reduced to 0.0001.
+This suggests that there exists an optimal (around 0.01) that best balances the two loss terms for optimal mIoU.
+The right plot explores the effect of LiDAR sweeps, referring to the number of point cloud frames aggregated over time.
+With sweeps increasing from 1 to 9, mIoU climbs steadily: starting at 50.91 (1 sweep), reaching 52.09 (3 sweeps), maintaining stability up to 5 sweeps, and then rising to 52.82 at 9 sweeps.
+Incorporating more sweeps generally enriches features from multi-frame data and improves mIoU performance, although the marginal gains diminish beyond a certain point.
+IV-D4 Ablation for Achieving Final Performance
+Table VI presents the ablation experiments that contributed to the final performance of DAOcc.
+First, extending the training time from 15 epochs to 24 epochs yields an improvement of 0.57 mIoU.
+Next, using CBGS [69] yields an additional 0.43 mIoU improvement.
+Since CBGS employs resampling, which increases the number of samples per epoch, we limit the training to 6 epochs to maintain a total training time comparable to that of the 24 epochs without CBGS for a fair comparison.
+Finally, increasing the learning rate can further improve performance.
+IV-D5 Comparison on Occ3D-Waymo with camera mask
+We further evaluate the robustness and generalization capability of DAOcc on the Occ3D-Waymo dataset, which remains a relatively underexplored benchmark for occupancy prediction.
+Following prior works [15, 65, 66], we train DAOcc using only 20% of the training data and discard LiDAR sweeps by utilizing a single LiDAR frame for each input, ensuring a fair comparison.
+As presented in Table VII, DAOcc significantly outperforms all existing methods, surpassing the multi-modal baseline BEVFormer-Fusion by 5.64 mIoU.
+Compared to the nuScenes dataset, which features a single top-mounted LiDAR, the Waymo dataset provides denser point clouds captured from five LiDARs.
+This richer sensory input further amplifies the performance gap between multi-modal and image-only methods.
+For instance, the recently proposed image-based method ODG [66] achieves an mIoU that is only 47.8% of that achieved by DAOcc (21.35 vs.
+44.69).
+These results highlight the critical importance of high-quality point cloud data in improving 3D occupancy prediction, particularly in complex environments, while also validating the strong generalization ability of DAOcc across different datasets and sensing setups.
+IV-D6 Robustness
+We aim to apply state-of-the-art 3D occupancy prediction models to real-world scenarios by leveraging sparse LiDAR data more effectively within a multi-modal fusion framework.
+Although robust alignment and fusion of sensor modalities are not the primary focus of this work, we nonetheless evaluate DAOcc’s robustness under challenging scenarios, as shown in Figure 7.
+In the nighttime scenario, under low illumination, the vehicle ahead (highlighted by the yellow dashed box) is barely observable in the camera image except for its lights.
+Even so, DAOcc successfully recognizes this vehicle and reconstructs a voxelized occupancy shape that closely matches the ground truth, demonstrating strong robustness in low-light environments.
+In the rainy scenario, raindrops obscure the camera view.
+The vehicle parked on the left (yellow dashed box) is nearly invisible, yet DAOcc still recognizes it accurately.
+A pedestrian partially occluded by another vehicle on the right (yellow dashed circle) is also correctly reconstructed in voxel space.
+Although DAOcc misses the vehicle immediately behind this pedestrian, this target is almost indiscernible even by manual inspection, and the corresponding LiDAR data also does not show vehicle traces at this position.
+We thus attribute this miss to the intrinsic difficulty of the sample rather than a failure of our method.
+The above results indicate that although DAOcc was not specifically designed for robustness, it already exhibits inherent robustness against interference due to the intrinsic generalization of data-driven learning paradigms.
+Based on this observation, we believe that increasing training data and designing specific data augmentation strategies for simulating sensor failures during training would further enhance DAOcc’s robustness.
+IV-D7 Detection Performance
+Since our proposed DAOcc incorporates an auxiliary 3D detection branch into the occupancy prediction framework, we also evaluate the performance of this auxiliary branch.
+The auxiliary detection branch identifies 10 object categories, including cars, trucks, buses, trailers, construction vehicles, pedestrians, motorcycles, bicycles, traffic cones, and barriers.
+We train the auxiliary 3D detection branch on the nuScenes dataset [24], and the results are evaluated using the mean average precision (mAP) and nuScenes detection score (NDS) metrics provided by the nuScenes benchmark.
+DAOcc achieves 59.4 mAP and 64.3 NDS on the nuScenes validation set, and the APs for the 10 object categories (in the order listed above) are 85.8, 55.5, 66.1, 30.8, 22.0, 75.8, 67.4, 52.2, 69.6, 69.2.
+These results show that DAOcc has the potential to be extended into a multi-task 3D perception framework.
+IV-E Visualization
 Figure 8 presents the visualizations of DAOcc on Occ3D-nuScenes validation set.
@@
 The results demonstrate that DAOcc is capable of making a relatively complete prediction of the scene, accurately reproducing it in fine detail.
-Dataset | Camera Mask | mIoU | mAP | NDS | Car | Truck | Bus | T.L. | C.V. | Ped. | M.C. | B.C. | T.C. | B.R. |
-Occ3D-nuScenes [44] | ✓ | 53.8 | 59.4 | 64.3 | 85.7 | 57.2 | 65.8 | 33.1 | 22.8 | 75.5 | 68.3 | 47.3 | 70.2 | 68.2 |
-Occ3D-nuScenes [44] | ✗ | 47.9 | 59.6 | 64.2 | 85.4 | 55.0 | 64.6 | 34.1 | 23.1 | 76.3 | 67.0 | 50.7 | 70.8 | 69.0 |
-SurroundOcc [51] | - | 30.5 | 53.6 | 60.2 | 81.9 | 42.7 | 56.0 | 26.3 | 19.7 | 72.6 | 57.7 | 43.8 | 66.7 | 68.1 |
+V Conclusion
+In this paper, we propose a novel multi-modal occupancy prediction framework, DAOcc, which aims to achieve superior performance while maintaining a deployment-friendly architecture with practical image backbones and input image resolutions.
+DAOcc achieves superior performance by fully leveraging the inherent geometry advantages of LiDAR data within a multi-modal framework, advantages largely overlooked by previous work.
+Extensive experiments on the Occ3D-nuScenes and Occ3D-Waymo datasets demonstrate that DAOcc surpasses existing methods in both accuracy and inference speed.
+Furthermore, due to its deployment-friendly design, DAOcc can be efficiently deployed and evaluated on edge devices.
+Notably, DAOcc runs in real-time on an AGX Orin device while maintaining state-of-the-art mIoU performance.